### v1

# Exact holomorphic solution of the Hodge-dual minimal surface: a confining quantum QCD string

Abstract
We present an exact analytic solution for the quaternionic Hodge–dual minimal surface in four dimensions, which provides a geometric fixed point of the Yang-Mills Wilson loop equation. The area of this surface is expressed through the eigenvalues of the Douglas–Gram matrix of the boundary loop, and is identical for the self–dual and anti–self–dual chiralities related by parity. We prove the Confinement Theorem by showing that the effective string tension $\sigma_{\mathrm{eff}}[C]$ (ratio of our Hodge dual area to a Euclidean one bounded by the same loop) is bounded as $\sqrt{2}\leq\sigma_{\mathrm{eff}}[C]\leq 2\sqrt{2}$, remaining finite and positive for all regular loops. Consequently, the Wilson loop decays exponentially with the Euclidean area, providing a purely geometric proof of confinement. Quantization of the quadratic holomorphic action yields a one-loop exact quantum string, with a universal $\tfrac{1}{2}\ln(2\pi)$ correction for each of the two chiral strings. The worldsheet theory is one-loop exact; the quantum corrections vanish after factoring out the self-dual rotational gauge orbit, leaving a purely spectral, gauge-invariant action. This construction defines a solvable quantum Hodge–dual string theory reproducing the QCD area law with explicit tension dependence on the loop shape.
[…]
[aff1]organization=Institute for Advanced Study, city=Princeton, country=USA
1 Introduction
Since the pioneering work of Ken Wilson [1], understanding confinement in Yang–Mills theory remains one of the central challenges of modern physics. Conventional approaches based on the local action obscure the geometric structure of the Wilson loop, where confinement is expected to emerge as an area law. In this work, we use the loop–space calculus [2] to reformulate the problem in purely geometric terms, leading to its analytic solution. In [2], the general Euler-Lagrange (EL) equations for the SD/ASD surface were derived, and an exact solution was found for the class of planar loops. We construct a Hodge-dual minimal surface, such that the corresponding area law exactly solves the loop equation for the Yang–Mills gradient flow. Here, we generalize this by presenting a complete holomorphic solution for arbitrary regular loops. This holomorphic solution supersedes the planar case constructed in Ref. [2], extending it to arbitrary loops and simultaneously correcting a typographical error $Y^{A}_{\mu}=0$ in the proof of the planar solution in that paper.
2 Basic definitions and duality equations
Our main goal is to minimize the self–dual (or anti–self–dual) area functional
$S_{\chi}[C;X]$ $=\min_{X\in\mathcal{M}_{\rm SD/ASD}}\int_{\mathbb{D}}\!\mathrm{d}^{2}z\,\sqrt{\tfrac{1}{2}\,\Sigma_{\mu\nu}\Sigma_{\mu\nu}},$ (2.1)
where $\Sigma_{\mu\nu}$ is the two–form constructed from the embedding $X_{\mu}(z,\bar{z})$ and defined below. The minimization is taken over embeddings $X_{\mu}(z,\bar{z})\in\mathbb{R}^{4}\!\otimes\!\mathbb{R}^{4}$ satisfying the self–duality (SD) or anti–self-duality (ASD) constraint $\ast\Sigma_{\mu\nu}=\pm\,\Sigma_{\mu\nu}$.
Quaternionic matrix structure. Let $M_{2}(\mathbb{C})$ carry the Hilbert–Schmidt inner product $\langle M,N\rangle_{\mathrm{HS}}=\tfrac{1}{2}\,\operatorname{tr}(MN^{\dagger})$ for $M,N\in M_{2}(\mathbb{C})$. Choose the orthonormal basis $T^{4}=I_{2}$, $T^{i}=\mathrm{i}\sigma_{i}$ ($i=1,2,3$), so that $\operatorname{tr}(T^{A}T^{B\dagger})=2\,\delta^{AB}$. For $X_{\mu}(z,\bar{z})\in M_{2}(\mathbb{C})$, $\mu=1,\dots,4$, define $X^{A}_{\mu}=\tfrac{1}{2}\,\operatorname{tr}(X_{\mu}T^{A\dagger})$, $X_{\mu}=\sum_{A=1}^{4}X^{A}_{\mu}\,T^{A}$.
Complexification of Cartesian components. $X^{A}_{z_{1}}=X^{A}_{1}+\mathrm{i}X^{A}_{2}$, $X^{A}_{\bar{z}_{1}}=X^{A}_{1}-\mathrm{i}X^{A}_{2}$; $X^{A}_{z_{2}}=X^{A}_{3}+\mathrm{i}X^{A}_{4}$, $X^{A}_{\bar{z}_{2}}=X^{A}_{3}-\mathrm{i}X^{A}_{4}$.
Two–form and Hodge decomposition. Define (with Einstein summation over $A$ implied)
$\Sigma_{\mu\nu}=\mathrm{i}\left(\partial_{z}X^{A}_{\mu}\,\partial_{\bar{z}}X^{A}_{\nu}-\partial_{\bar{z}}X^{A}_{\mu}\,\partial_{z}X^{A}_{\nu}\right).$ (2.2)
With $(\ast\Sigma)_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\rho\sigma}\Sigma_{\rho\sigma}$, the SD/ASD parts are $\Sigma^{(\pm)}_{\mu\nu}=\tfrac{1}{2}(\Sigma_{\mu\nu}\pm\ast\Sigma_{\mu\nu})$, satisfying $\ast\Sigma^{(\pm)}_{\mu\nu}=\pm\Sigma^{(\pm)}_{\mu\nu}$. The Euler-Lagrange variational equations for this functional were written down in [2]. These nonlinear differential equations were solved in that paper only for a special case of planar boundary loop $C$. In this paper, we generalize that solution for an arbitrary loop.
Self-dual area derivative. There is an internal $O(4)$ gauge invariance of our area element in addition to a physical $O(4)$ invariance of the problem in $\mathbb{R}^{4}$. We use that extra degree of freedom to provide the self-duality (SD or ASD) of the area derivative [2]. The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
$\frac{\delta W}{\delta\sigma_{\mu\nu}(\theta)}=$
$\frac{\delta}{\delta C^{\prime}_{\mu}(\theta-0)}\frac{\delta W}{\delta C^{\prime}_{\nu}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\mu}(\theta+0)}\frac{\delta W}{\delta C^{\prime}_{\nu}(\theta-0)}$ (2.3)
The loop equation for the Yang–Mills gradient flow was shown in [2] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
$\mathcal{L}(S)=0;$ (2.4)
$\mathcal{L}=\oint_{\theta}C^{\prime}_{\nu}(\theta)\left(\frac{\delta}{\delta C^{\prime}_{\mu}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\mu}(\theta-0)}\right)\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}$ (2.5)
The basic observation made in that paper was there is a SD solution similar to the Yang–Mills multi-instanton. The loop space diffusion equation (2.5) is identically satisfied for any functional $S[C]$ with SD or ASD area derivative
$\ast\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)}=\pm\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)};$ (2.6)
$e_{\alpha\mu\nu\lambda}\left(\frac{\delta}{\delta C^{\prime}_{\alpha}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\alpha}(\theta-0)}\right)\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)}\equiv 0;$ (2.7)
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [2] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents SD or ASD tensor. This would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
3 Geometric Definition of the Physical Area
The construction of the action $S_{\chi}$ in Sec. 4 depends on a specific choice of boundary conditions (BCs), which introduces a rotational ambiguity. The resulting functional is not $O(4)$ invariant. We must therefore provide a geometric definition of the physical area $S_{\mathrm{phys}}[C]$ that is fully invariant.
We define $S_{\mathrm{phys}}[C]$ as the solution of the minimax problem: first, we minimize the value of the Hodge-dual action $S[X]$ (2.1), by solving the Euler-Lagrange (E-L) equations for $S$ with specific boundary conditions $X=Q.C$ using a normalized matrix $Q$ compatible with the self-duality of the area derivative (see below). Then, we maximize this minimal value over the full group of allowed isometries of $Q$ (transformations that leave the E-L equations invariant).
Case 1: Simple (non-intersecting) loops
For a single, non-intersecting loop $C$, the E-L equations and local Dirichlet BCs $X=QC$ are global. The only relevant isometry that preserves the problem is a global $O(4)$ rotation, $R$. The rotation of $Q$ is equivalent to the opposite rotation of $C$ by the invariance of our area integral. The set of compatible solutions is thus $\{X[RC]\mid R\in O(4)\}$. The definition of the area becomes:
$S_{\mathrm{phys}}[C]:=\max_{R\in O(4)}S[X[RC]]$ (3.1)
This is precisely the definition that will be evaluated in Sec. 5, which yields the $O(4)$-invariant spectral law.
Case 2: Composite (self-intersecting) loops
For a loop that is a composite of two (or more) separated or self-intersecting sub-loops, $C=C_{1}\uplus C_{2}$, the problem changes. The E-L equations and the Dirichlet BCs are local. The problem of finding the minimal surface $X$ for the boundary $C$ therefore decouples into two independent minimal surface problems: $X_{1}$ on domain $\mathbb{D}_{1}$ bounded by $Q_{1}C_{1}$, and $X_{2}$ on domain $\mathbb{D}_{2}$ bounded by $Q_{2}C_{2}$. The total action for the solution is the sum $S[X]=S[X_{1}]+S[X_{2}]$.
Because the problem decouples, the isometry group of the total solution is larger. It is not the "diagonal" $O(4)$ group (which would force $R_{1}=R_{2}$), but the full product group $O(4)_{1}\times O(4)_{2}$. A transformation $(R_{1},R_{2})$ with $R_{1}\neq R_{2}$ is a valid isometry of the decoupled E-L equations.
To find the true maximum over "all compatible isometries" (per our definition), we must maximize over this larger group. This defines the area for a composite loop as $S_{\mathrm{phys}}[C_{1}\uplus C_{2}]=\max_{R_{1}\in O(4)}S[X_{1}[R_{1}C_{1}]]+\max_{R_{2}\in O(4)}S[X_{2}[R_{2}C_{2}]]$.
4 A Self-Dual Holomorphic Ansatz
Holomorphic map.
We work in the target complex basis $z_{1}=x_{1}+\mathrm{i}x_{2}$, $z_{2}=x_{3}+\mathrm{i}x_{4}$. Let $U^{A}_{z_{a}}(z)$ be the holomorphic $(1,0)$ components of the embedding $X^{A}_{\mu}(z,\bar{z})=U^{A}_{\mu}(z)+\overline{U^{A}_{\mu}(z)}$. We also define $V^{A}_{z_{a}}:=\partial_{z}U^{A}_{z_{a}}$. As we shall see, the EL equations for the minimal Hodge-dual surface are identically satisfied by virtue of the Cauchy-Riemann equations for the holomorphic functions $U^{A}_{\mu}(z)$ combined with the SD property of the area element.
Lemma 4.1 (Normalization).
With the canonical SD $(1,1)$ form $B:=\mathrm{i}\big(dz_{1}\!\wedge d\bar{z}_{1}+dz_{2}\!\wedge d\bar{z}_{2}\big)$, the Euclidean Hodge star gives $\sqrt{\tfrac{1}{2}\,B_{\mu\nu}B_{\mu\nu}}=2\sqrt{2}$. Consequently, the area density of any SD surface $\Sigma=a(z)B$ is
$\sqrt{\tfrac{1}{2}\,\Sigma^{2}}\;$ = $=\;2\sqrt{2}\;a(z)$
$=\;\sqrt{2}\sum_{A}\big(|V^{A}_{z_{1}}|^{2}+|V^{A}_{z_{2}}|^{2}\big),$ (4.1)
where the second line follows from $\Sigma_{z_{1}\bar{z}_{1}}=\Sigma_{z_{2}\bar{z}_{2}}=\mathrm{i}a(z)$ and the definition $\Sigma_{z_{a}\bar{z}_{b}}=\mathrm{i}\sum_{A}V^{A}_{z_{a}}\overline{V^{A}_{z_{b}}}$.
Proof.
In the real coframe $e^{1}=dx_{1},\dots,e^{4}=dx_{4}$, $\mathrm{i}dz_{1}\!\wedge d\bar{z}_{1}=2\,e^{1}\!\wedge e^{2}$ and $\mathrm{i}dz_{2}\!\wedge d\bar{z}_{2}=2\,e^{3}\!\wedge e^{4}$, so $B=2(e^{1}\!\wedge e^{2}+e^{3}\!\wedge e^{4})$. Then $\frac{1}{2}B^{2}=8$ and the identities follow. ∎
Two–form and Self-Duality.
The two-form is $\Sigma=\mathrm{d}U\wedge\mathrm{d}\overline{U}$. As $U$ is holomorphic, $\Sigma$ is of pure $(1,1)$ type. The self-duality condition $\ast\Sigma=\Sigma$ requires $\Sigma_{z_{1}\bar{z}_{1}}=\Sigma_{z_{2}\bar{z}_{2}}$. Using the component definition $\Sigma_{z_{a}\bar{z}_{b}}=\mathrm{i}\sum_{A}V^{A}_{z_{a}}\overline{V^{A}_{z_{b}}}$ from Lemma 4.1, this condition implies the explicit constraint:
$\sum_{A}\big(|V^{A}_{z_{1}}|^{2}\big)=\sum_{A}\big(|V^{A}_{z_{2}}|^{2}\big).$ (4.2)
The area density is given by Eq. (4.1).
An Explicit SD Ansatz.
We propose a minimal ansatz that satisfies the self-duality condition (4.2) using two independent holomorphic functions $f_{1}(z)$ and $f_{2}(z)$. Let the only non-zero components of $U$ be:
$U^{4}_{z_{1}}=f_{1}(z),\hskip 18.49988ptU^{3}_{z_{2}}=f_{1}(z),$
$U^{2}_{z_{1}}=f_{2}(z),\hskip 18.49988ptU^{1}_{z_{2}}=f_{2}(z).$ (4.3)
This choice manifestly satisfies (4.2), as $\sum_{A}|V^{A}_{z_{1}}|^{2}=|f_{1}^{\prime}|^{2}+|f_{2}^{\prime}|^{2}$ and $\sum_{A}|V^{A}_{z_{2}}|^{2}=|f_{2}^{\prime}|^{2}+|f_{1}^{\prime}|^{2}$.
Boundary Conditions and Explicit Projector.
The action functional (4.1) for this ansatz becomes a Dirichlet functional
$S_{\chi}=2\sqrt{2}\int_{\mathbb{D}}\big(|f_{1}^{\prime}|^{2}+|f_{2}^{\prime}|^{2}\big)\mathrm{d}^{2}z.$ (4.4)
The EL equations of this functional are satisfied by Cauchy-Riemann equations for $f_{1},f_{2}$, so that this functional reduces to the boundary data. By the Dirichlet identity, $S_{\chi}=2\sqrt{2}\sum_{n\geq 1}n\big(|\hat{f}_{1,n}|^{2}+|\hat{f}_{2,n}|^{2}\big)$. We now fix the boundary conditions by relating $f_{1},f_{2}$ to the boundary loop $C_{\mu}$ in a rotated basis:
$\Re f_{1}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{1}(\theta)+C_{2}(\theta)\big),$
$\Re f_{2}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{3}(\theta)+C_{4}(\theta)\big).$ (4.5)
Substituting the Fourier modes $\hat{f}_{1,n}=\frac{1}{\sqrt{2}}(\hat{C}_{n,1}+\hat{C}_{n,2})$ and $\hat{f}_{2,n}=\frac{1}{\sqrt{2}}(\hat{C}_{n,3}+\hat{C}_{n,4})$ into (4.4):
$S_{\chi}$ $=2\sqrt{2}\sum_{n\geq 1}n\left(\Big|\frac{\hat{C}_{n,1}+\hat{C}_{n,2}}{\sqrt{2}}\Big|^{2}+\Big|\frac{\hat{C}_{n,3}+\hat{C}_{n,4}}{\sqrt{2}}\Big|^{2}\right)$
$=\sqrt{2}\sum_{n\geq 1}n\left(|\hat{C}_{n,1}+\hat{C}_{n,2}|^{2}+|\hat{C}_{n,3}+\hat{C}_{n,4}|^{2}\right).$ (4.6)
Introducing the Douglas-Gram matrix $G_{\mu\nu}:=\sum_{n\geq 1}n\,\hat{C}_{n,\mu}\,\overline{\hat{C}_{n,\nu}}$, this action is a quadratic form $S_{\chi}=2\sqrt{2}\,\operatorname{tr}(PG)$, where $P$ is the explicit rank-2 projector $P=\frac{1}{2}\left(\begin{smallmatrix}1&1&0&0\\ 1&1&0&0\\ 0&0&1&1\\ 0&0&1&1\end{smallmatrix}\right)$.
Lemma 4.2 (Projector Property).
The matrix $P$ defined by (4) is a rank-2 orthogonal projector ($P\in\operatorname{Gr}(2,4)$). This is verified by direct computation: $P$ is symmetric, $P^{2}=P$, and $\operatorname{tr}P=2$, so $\operatorname{rank}(P)=2$.
Admissible loops.
Throughout we assume the boundary loop $C:S^{1}\to\mathbb{R}^{4}$ is admissible in the sense that the spectral expansion converges in the Douglas–Gram matrix, so that the Dirichlet identity leading to (4.6) is valid, and that the Fourier/Hilbert-transform manipulations used in this paper are justified.
5 Invariant Area and the Spectral Law
Rotationally invariant area.
The true, physical $O(4)$-invariant area $S_{\mathrm{phys}}$ was defined as the maximum area that can be generated by our functional form over all possible orientations $R$ of the loop $C$. For this holomorphic solution
$S_{\mathrm{phys}}[\mathcal{C}]$ $\;:=\;\max_{R\in O(4)}S_{\chi}[RC;P]$
$\;=\;\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}\!\left(P\,(RGR^{\top})\right).$ (5.1)
Let $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$ be the eigenvalues of $G$. The maximum in (5.1) is evaluated by first applying the cyclic property of the trace, which gives $\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}((R^{T}PR)G)$. We can define a new projector $P^{\prime}=R^{T}PR$. Since $P$ is a fixed rank-2 projector, $P^{\prime}$ is also a rank-2 projector. The group $O(4)$ acts transitively on the space of all 2-planes, $\operatorname{Gr}(2,4)$. This means that by maximizing over all $R\in O(4)$, the projector $P^{\prime}$ sweeps through the entire set of all possible rank-2 projectors. Therefore, our maximization is mathematically identical to $2\sqrt{2}\max_{P^{\prime}\in\operatorname{Gr}(2,4)}\operatorname{tr}(P^{\prime}G)$. By the Ky Fan theorem [3], this maximum is precisely the sum of the two largest eigenvalues, $\lambda_{3}+\lambda_{4}$.
Outcome and the spectral law. Crucially, the result of this maximization is $S_{\mathrm{phys}}[\mathcal{C}]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})$. This value is a universal spectral invariant, independent of the specific rank-2 projector $P$ we started with.
$S_{\mathrm{phys}}[\mathcal{C}]\;=\;2\sqrt{2}\,(\lambda_{3}+\lambda_{4}).$ (5.2)
This calculation confirms that our physical definition of the invariant area (5.1) leads directly to the $O(4)$-invariant spectral law. For planar loops ($\operatorname{rank}G=2$), $\lambda_{1}=\lambda_{2}=0$, and this correctly specializes to $S_{\mathrm{phys}}[\mathcal{C}]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})=2\sqrt{2}\,\mathcal{D}[C]$, reproducing the claim of Theorem 1 of [2]. The optimal projector $P_{\ast}=v_{3}v_{3}^{\top}+v_{4}v_{4}^{\top}$ (where $v_{i}$ are eigenvectors of $G$) is the one that achieves this maximum.
6 Self-Duality of the Area Derivative
Theorem 6.1 (Self-Duality of the Physical Action).
The physical action $S_{\mathrm{phys}}[\mathcal{C}]$ defined in (5.2) has a self-dual (SD) area derivative:
$\ast\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}=\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}.$ (6.1)
Proof.
The physical action $S_{\mathrm{phys}}[\mathcal{C}]$ is the envelope (the maximum) of a family of functionals $S_{R}[C]:=S_{\chi}[RC;P]$, indexed by the rotation $R\in O(4)$, where $S_{\chi}[C;P]$ is the base functional from Sec. 4. The area derivative operator $\mathcal{D}:=\frac{\delta}{\delta\sigma}$ and the Hodge star $\ast$ are both $O(4)$-covariant. This means that if we can prove self-duality for any single member of this family (e.g., the $R=I$ case, $S_{\chi}[C;P]$), then self-duality holds for all members. The derivative of the envelope $S_{\mathrm{phys}}$ at a loop $C$ is, by definition, the derivative of the specific functional $S_{R_{\text{max}}}[C]$ that achieves the maximum at that point. Since all functionals in the family will be shown to have an SD derivative, $S_{\mathrm{phys}}$ must also have an SD derivative. Therefore, the proof reduces to verifying the self-duality for the base functional $S_{\chi}[C;P]$ (for $R=I$), which we do now.
We begin with the generic formula from [2] for the first variation of the Hodge-dual area
$\delta S=2\int d\theta\delta X^{A}_{\mu}(\theta)\partial_{\theta}X^{A}_{\nu}(\theta)T_{\mu\nu}$ (6.2)
where $T_{\mu\nu}$ is the normalized area element $\Sigma_{\mu\nu}$. [^1] This is generic variation of our functional in the space of all functions $X^{A}_{\mu}(u,v)$, evaluated on our holomorphic solution after taking a variation. Had we computed the variation on the class of holomorphic functions by reducing the Dirichlet functional to the boundary, we would have found a symmetric quadratic form with a vanishing (and trivial) area derivative. Thus, variations orthogonal to the solution manifold are essential for the non-trivial area derivative. With our holomorphic solution, using the constant $B_{\mu\nu}$ tensor and the Dirichlet functional with the Hilbert transform, we find
$\delta S\propto\iint d\theta_{1}d\theta_{2}H(\theta_{1}-\theta_{2})Q(C(\theta_{1}),\delta C^{\prime}(\theta_{2}));$ (6.3)
$Q(A,B)=A_{2}B_{1}-A_{1}B_{2}+A_{3}B_{4}-A_{4}B_{3}$ (6.4)
Now, the first and second variations are
$\frac{\delta S}{\delta C^{\prime}(\theta_{2})}\propto\int d\theta_{1}H(\theta_{1}-\theta_{2})Q(C(\theta_{1}),I_{4});$ (6.5)
$\left.\frac{\delta^{2}S}{\delta C^{\prime}_{\mu}(\theta^{\prime})\delta C^{\prime}_{\nu}(\theta)}\right|_{\theta-0}^{\theta+0}\propto\int_{0}^{\theta}d\theta^{\prime}H(\theta^{\prime}-\theta)\hat{Q}_{\mu\nu};$ (6.6)
where $\hat{Q}_{\mu\nu}$ is the constant, antisymmetric tensor $\hat{Q}=\text{diag}(i\sigma_{2},i\sigma_{2})$. In the second derivative, by the chain rule, we use the standard Volterra derivative [4] of the loop with respect to its tangent, $\delta C(x)/\delta C^{\prime}(y)=\Theta(x-y)$. This tensor $\hat{Q}$ is manifestly SD (as $\ast\hat{Q}_{12}=\hat{Q}_{34}$ and $\ast\hat{Q}_{34}=\hat{Q}_{12}$), which proves the area derivative for $S_{\chi}[C;P]$ is self-dual. By the covariance argument above, this property extends to $S_{\mathrm{phys}}$.
Note that the numerical factor $\int_{0}^{\theta}d\theta^{\prime}H(\theta^{\prime}-\theta)\propto\sum_{n>0}1/n$ logarithmically diverges. However, the SD of the area derivative, valid for the tensor factor $Q_{\mu\nu}$, leads to the cancelation of the whole series in the loop space diffusion equation (2.5) (as argued in Sec. 10). Note that this cancelation is valid irrespective of the cutoff at large $n$ in Fourier series, so that if we fix this cutoff and substitute the area derivative into the loop equation (2.5), we get exactly zero. Thus, we can still claim the finite solution of the loop equation , as the minimal area is finite, and its contribution to the loop space diffusion equation is zero, making it a fixed point. ∎
7 Reparametrization and additivity
Parametric invariance. As stated below (5.1) and in Sec. 3, the geometric area is defined by maximization over allowed isometries; we additionally take the infimum over orientation–preserving reparametrizations to remove any dependence on the boundary parameter. We therefore adopt the parametrization–invariant form
$S_{\mathrm{phys}}[C]$
$:=\min_{\phi\in\mathrm{Diff}^{+}(S^{1})}\;\max_{R\in O(4)}\,2\sqrt{2}\;\operatorname{tr}\!\Big(P\,R\,G[C\!\circ\!\phi]\,R^{\top}\Big).$ (7.1)
By construction, the choice of parameter is without loss of generality: one may first reparametrize to a convenient standard $0,2\pi]$ and only then perform the $O(4)$ maximization and the outer minimum over $\phi$.
Canonical split at a self–intersection. Let $C(\theta_{1})=C(\theta_{2})$ be a transversal self–intersection and define the two oriented subloops $C_{1}=C[\theta_{1},\theta_{2}]$ and $C_{2}=C[\theta_{2},\theta_{1}\!+\!2\pi]$ as in Sec. 3 (Case 2). Using parametric invariance, choose $\phi_{1},\phi_{2}\in\mathrm{Diff}^{+}(S^{1})$ with $\phi_{1}(S^{1})=[\theta_{1},\theta_{2}]$ and $\phi_{2}(S^{1})=[\theta_{2},\theta_{1}\!+\!2\pi]$ and set $\widetilde{C}_{j}(\tau)=C(\phi_{j}(\tau))$ so that each boundary component is on $0,2\pi]$. For a disconnected boundary the Douglas–Gram series splits,
$G[\widetilde{C}_{1}\!\uplus\!\widetilde{C}_{2}]$ $=G[\widetilde{C}_{1}]+G[\widetilde{C}_{2}],$ (7.2)
and, as explained in Sec. 3 (Case 2), the E–L problems decouple and the isometry group enlarges to the product $O(4)_{1}\times O(4)_{2}$.
Additivity of $S_{\mathrm{phys}}$. With (7.2), the double extremization factorizes:
$S_{\mathrm{phys}}[C_{1}\!\uplus\!C_{2}]:=2\sqrt{2}\,$
$\min_{\phi_{1},\phi_{2}}\;\max_{\begin{subarray}{c}R_{1}\in O(4)\\ R_{2}\in O(4)\end{subarray}}\Big\{\operatorname{tr}\!\big(P\,R_{1}G[C_{1}\!\circ\!\phi_{1}]R_{1}^{\top}\big)$
+ $\hskip 73.07445pt+\,2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{2}G[C_{2}\!\circ\!\phi_{2}]R_{2}^{\top}\big)\Big\}$ (7.3)
$=\min_{\phi_{1}}\max_{R_{1}}2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{1}G[C_{1}\!\circ\!\phi_{1}]R_{1}^{\top}\big)$
$+\;\min_{\phi_{2}}\max_{R_{2}}2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{2}G[C_{2}\!\circ\!\phi_{2}]R_{2}^{\top}\big)$
$=\;S_{\mathrm{phys}}[C_{1}]+S_{\mathrm{phys}}[C_{2}].$ (7.4)
The first equality is the definition consistent with Sec. 3 (Case 2); the second uses the product structure of the objective (sum of two independent functionals), which makes the $\max$ over $R_{1},R_{2}$ and the $\min$ over $\phi_{1},\phi_{2}$ separate.
Compatibility with the MM kernel. The right–hand side of the MM equation [5, 6] is supported on self–intersections via the bilocal kernel $\int\!\!\int d\theta_{1}d\theta_{2}\,C^{\prime}(\theta_{1})\!\cdot\!C^{\prime}(\theta_{2})\,\delta(C(\theta_{1})-C(\theta_{2}))\,\cdots$. Under $\theta_{a}=\phi_{a}(\tau_{a})$ (with $\phi_{a}^{\prime}\!>\!0$),
$d\theta_{a}\,C^{\prime}(\theta_{a})$ $=d\tau_{a}\,\dot{C}(\tau_{a}),$
$\delta\!\big(C(\theta_{1})-C(\theta_{2})\big)$ $=\delta\!\big(C(\phi_{1}(\tau_{1}))-C(\phi_{2}(\tau_{2}))\big),$ (7.5)
so the kernel is parametrization–invariant. Therefore the canonical standardization to $0,2\pi]$ used above is legitimate and (7.4) guarantees that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\uplus\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.5) and [7, 2]).
8 Confinement theorem
Theorem 8.1 (Confinement area law).
Let $G$ be the Douglas–Gram matrix of $C$ with eigenvalues $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$. as we have proven in the previous section, the Hodge–dual area and the Euclidean minimal area satisfy
$A_{\mathrm{geom}}[C]:=\operatorname{tr}G\;=\;\sum_{i=1}^{4}\lambda_{i},\qquad S_{\chi}[C]\;=\;2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$ (8.1)
$\sigma_{\mathrm{eff}}[C]:=\frac{S_{\chi}[C]}{A_{\mathrm{geom}}[C]}\;=\;2\sqrt{2}\,\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}.$ (8.2)
Moreover,
[…]
with equality on the right for planar spectra ($\lambda_{1}=\lambda_{2}=0$),and on the left for isotropic spectra ($\lambda_{1}=\cdots=\lambda_{4}$).
The quantity $\sigma_{\mathrm{eff}}[C]$ (hence $S_{\chi}[C]$) is O(4)O(4)–invariant and chirality–even: if $C\mapsto RC$ with $R\in O(4)$, then $G\mapsto RGR^{\top}$ leaves $\{\lambda_{i}\}$, and thus the theorem, unchanged.
Remark 8.1 (Minimization over reparametrization).
Both functionals, Douglas’s $\sum\lambda_{i}$ and our $\lambda_{3}+\lambda_{4}$, have yet to be minimized over the reparametrization of the bounding loop $C(\theta)\Rightarrow C(f(\theta)),f^{\prime}>0$. As the minimization of just the two highest eigenvalues, compared to that of the trace, is attained at different parametrizations, we can only state that $\sigma_{eff}=2\sqrt{2}\frac{\min(\lambda_{3}+\lambda_{4})}{\min(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})}$. However, the lower inequality still holds $\sigma_{eff}\geq\sqrt{2}$, as for every parametrization $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}\leq 2(\lambda_{3}+\lambda_{4})$.Therefore, the minimum in the denominator is less than or equal to twice that in the numerator, proving the lower bound quoted in confinement theorem.
9 Quantization
Our area functional is equivalent to the string theory with action $S_{0}[f_{i}]=\int_{\mathbb{D}}\!\mathrm{d}^{2}z\,(\nabla f_{i})^{2}+\int_{\mathbb{D}}\!\mathrm{d}^{2}z\,u_{i}\,\partial_{\bar{z}}f_{i}$, where $f_{1,2}(z)$ are the basis functions introduced in Section 4. On the boundary, we have linear boundary conditions of the form: $S_{\partial}[f_{i},\rho]=\oint\!\mathrm{d}\theta\,\rho_{i}(\theta)[\Re f_{i}\left(e^{\mathrm{i}\theta}\right)-Q_{i}(C(\theta))]$ with $Q_{1}(C)=(C_{1}+C_{2})/\sqrt{2},Q_{2}(C)=(C_{3}+C_{4})/\sqrt{2}$. Integrating out $f_{i}$ reproduces a Douglas-type boundary action $\mathcal{D}[QC]=\operatorname{tr}PG$. The functional integral over $f_{i}$ defines the quantum Hodge–dual string. The area derivative’s self-duality stems from local Hodge duality and holds quantum-mechanically. The quadratic action ensures quantum Schwinger–Dyson equations match classical E-L relations. Thus, the expectation value $\left\langle\delta S/\delta\sigma\right\rangle$ is self-dual, and the quantum string obeys the classical loop equations. Integrating out $f_{i}$ gives a Gaussian integral $Z[C;P]\!\propto\![\det{}^{\prime}(K_{P})]^{-1/2}\exp\left(-\,2\sqrt{2}\,\kappa\operatorname{tr}PG\right)$, where $K_{P}=P\,H\,\partial_{\theta}P$ has a universal spectrum. Hence $\det{}^{\prime}(K_{P})$ is a constant, and the effective action is classical: $\Gamma[C;P]=2\sqrt{2}\,\operatorname{tr}PG+\text{const}$. Since $S_{\chi}[C;P]=2\sqrt{2}\,\operatorname{tr}(PG)$ is spectral, $SO(3)_{+}$ rotations of $P$ are a gauge orbit. The integration factors out, and $Z[C]$ is evaluated on the spectral projector $P_{\ast}=v_{3}v_{3}^{\top}+v_{4}v_{4}^{\top}$: $Z[C]\;\propto\;\exp\left(-\,2\sqrt{2}\,\kappa(\lambda_{3}+\lambda_{4})\right)$. The gauge orbit carries no dynamics; $P_{\ast}$ captures the full invariant content.
10 Application to QCD
The known UV corrections (perimeter/cusp terms) to the QCD Wilson loop are sub-leading at large loop size $L\to\infty$. As we argue in detail in [2], due to the Leibniz property of the loop diffusion operator $\mathcal{L}$, the transformation $W[C]\Rightarrow W[C]\exp\left(-\kappa S[C]\right)$ leaves the MM loop equation[7] $\mathcal{L}W[C_{1}\uplus C_{2}]=W[C_{1}]\otimes W[C_{2}]$ invariant if $\mathcal{L}S[C]=0$ and $S[C]$ are additive on self-intersecting loops $S[C_{1}\uplus C_{2}]=S[C_{1}]+S[C_{2}]$. Our SD/ASD area satisfies both of these conditions (Euclidean area satisfies only additivity). SD or ASD areas satisfy the equation $\mathcal{L}S=0$ by virtue of the loop space Bianchi identity [2], just as the SD Yang–Mills field satisfies the equations of motion. Given the perturbative solution $W_{0}[C]$ we have the general solution as $W[C]=W_{0}[C]\exp\left(-\kappa S[C]\right)$, where the string mass scale $\kappa$ must be fixed as a universal number times the QCD mass scale by matching the perturbative region of small loops.
Crucially, the MM equation requires factorization of the loop derivative $\mathcal{L}W$, not the functional itself. As shown in [7], while the perturbative $W_{0}$ retains non-trivial correlations between sub-loops, the graphs connecting these sub-loops cancel upon application of $\mathcal{L}$, leaving exactly the factorized contact term on the right hand side. Our ansatz remains consistent because the geometric action is strictly additive, $S[C_{1}\oplus C_{2}]=S[C_{1}]+S[C_{2}]$. Combined with the zero-mode property $\mathcal{L}S=0$, the factor $\exp\left(-\kappa S\right)$ cancels identically from both sides of the equation, allowing the confining factor to dress $W_{0}$ without constraining its internal correlations.
11 Conclusion
We have presented an exact solution for the quaternionic Hodge–dual minimal surface, a new class of minimal surfaces in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ that is exactly solvable by reduction to free conformal maps. The minimized area is the non-trivial spectral invariant $S_{\min}(C)=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})$, where $\{\lambda_{i}\}$ are the eigenvalues of the Douglas–Gram matrix $G$. This result is mathematically robust: the optimal surface is determined by the two dominant Douglas eigenvectors, and the underlying worldsheet theory is quantum-exact.
This Hodge–dual minimal area $S[C]$ enters the non-perturbative factor $\exp\left(-\kappa S[C]\right)$ multiplying the perturbative Wilson loop, solving the MM equation in QCD [5, 6, 7].
The primary physical implication is a first-principles geometric proof of confinement, as stated in our Confinement Theorem (Theorem 8.1). This theorem provides an explicit, shape-dependent formula for the effective string tension $\sigma_{\mathrm{eff}}[C]$ (defined as a ratio of $S[C]$ to the Euclidean area $A[C]$). This parameter depends on the loop shape and is bounded from below as $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$. This solution lays the groundwork for a quantitative theory of the high-level hadron spectrum.
Acknowledgements.—The author thanks colleagues at the Institute for Advanced Study for inspiring discussions.