### v2

# Geometric Solution of the Loop Equation and Quark Confinement in Planar QCD

Abstract
We present an exact analytic area law for planar QCD satisfying the MM loop equation for arbitrary loops. The solution is constructed using a quaternionic Hodge-dual minimal surface in a 16-dimensional auxiliary space. Unlike random surfaces, this construction defines a topological string theory where the action is proportional to the area of the conformal map of the 2D disk, possessing no bulk quantum fluctuations.
The Wilson loop factorizes into a confining dressing factor and the perturbative term: $W[C]=W_{pert}[C]e^{-\kappa S[C]}$. This corresponds to a non-perturbative gluon condensate vacuum supporting planar graph fluctuations. We prove the inequality $S[C]\geq\sqrt{2}A[C]$ (where $A[C]$ is the Euclidean minimal area), establishing a sufficient condition for confinement.
The area law is defined explicitly as $S[C]=2\sqrt{2}(\lambda_{3}+\lambda_{4})$ using the two largest eigenvalues of the Douglas-Gram matrix. We demonstrate that this functional is an exact additive zero mode of the loop diffusion operator, protected algebraically by the Bianchi identity. This ensures global validity for all regular loops ($C\in H^{1/2}(\mathbb{S}^{1},\mathbb{R}^{4})$).
Matching this confining factor with the gluon condensate via the QCD Operator Product Expansion (OPE) determines the physical string tension $\kappa\sim\Lambda_{QCD}^{2}$ (up to a universal numerical factor). This framework provides a first-principles derivation of confinement in large-$N_{c}$ QCD and new tools for the high-lying meson spectrum.
[…]
[aff1]organization=Institute for Advanced Study, city=Princeton, country=USA
1 Introduction
Since the pioneering work of Ken Wilson [Wilson1974Confinement], understanding confinement in Yang–Mills theory remains one of the central challenges of modern physics. Conventional approaches based on the local action obscure the geometric structure of the Wilson loop, where confinement is expected to emerge as an area law. In this work, we use the loop–space calculus [migdal2025SQYMflow] to reformulate the problem in purely geometric terms, leading to its analytical solution. In [migdal2025SQYMflow], the general Euler-Lagrange (EL) equations for the SD/ASD surface were derived, and an exact solution was found for the class of planar loops. We construct a Hodge-dual minimal surface, such that the corresponding area law exactly solves the loop equation for the Yang–Mills gradient flow. Here, we generalize this by presenting a complete holomorphic solution for arbitrary regular loops (with Fourier coefficients decreasing faster than $1/n$). This holomorphic solution supersedes the planar case constructed in Ref. [migdal2025SQYMflow], extending it to arbitrary loops and simultaneously correcting a typographical error $Y^{A}_{\mu}=0$ in the proof of the planar solution in that paper.
2 Basic definitions and duality equations
Our main goal is to minimize the self–dual (or anti–self–dual) area functional
$S_{\chi}[C;X]$ $=\min_{X\in\mathcal{M}_{\rm SD/ASD}}\int_{\mathcal{D}}\!\mathrm{d}^{2}z\,\sqrt{\tfrac{1}{2}\,\Sigma_{\mu\nu}\Sigma_{\mu\nu}},$ (2.1)
where $\Sigma_{\mu\nu}$ is the two–form constructed from the embedding $X_{\mu}(z,\bar{z})$ and defined below. The minimization is taken over embeddings $X_{\mu}(z,\bar{z})\in\mathbb{R}^{4}\!\otimes\!\mathbb{R}^{4}$ satisfying the self–duality (SD) or anti–self-duality (ASD) constraint $\ast\Sigma_{\mu\nu}=\pm\,\Sigma_{\mu\nu}$.
Quaternionic matrix structure. Let $M_{2}(\mathbb{C})$ carry the Hilbert–Schmidt inner product $\langle M,N\rangle_{\mathrm{HS}}=\tfrac{1}{2}\,\operatorname{tr}(MN^{\dagger})$ for $M,N\in M_{2}(\mathbb{C})$. Choose the orthonormal basis $T^{4}=I_{2}$, $T^{i}=\mathrm{i}\sigma_{i}$ ($i=1,2,3$), so that $\operatorname{tr}(T^{A}T^{B\dagger})=2\,\delta^{AB}$. For $X_{\mu}(z,\bar{z})\in M_{2}(\mathbb{C})$, $\mu=1,\dots,4$, define $X^{A}_{\mu}=\tfrac{1}{2}\,\operatorname{tr}(X_{\mu}T^{A\dagger})$, $X_{\mu}=\sum_{A=1}^{4}X^{A}_{\mu}\,T^{A}$.
Complexification of Cartesian components. $X^{A}_{z_{1}}=X^{A}_{1}+\mathrm{i}X^{A}_{2}$, $X^{A}_{\bar{z}_{1}}=X^{A}_{1}-\mathrm{i}X^{A}_{2}$; $X^{A}_{z_{2}}=X^{A}_{3}+\mathrm{i}X^{A}_{4}$, $X^{A}_{\bar{z}_{2}}=X^{A}_{3}-\mathrm{i}X^{A}_{4}$.
Two–form and Hodge decomposition. Define (with Einstein summation over $A$ implied)
$\Sigma_{\mu\nu}=\mathrm{i}\left(\partial_{z}X^{A}_{\mu}\,\partial_{\bar{z}}X^{A}_{\nu}-\partial_{\bar{z}}X^{A}_{\mu}\,\partial_{z}X^{A}_{\nu}\right).$ (2.2)
With $(\ast\Sigma)_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\rho\sigma}\Sigma_{\rho\sigma}$, the SD/ASD parts are $\Sigma^{(\pm)}_{\mu\nu}=\tfrac{1}{2}(\Sigma_{\mu\nu}\pm\ast\Sigma_{\mu\nu})$, satisfying $\ast\Sigma^{(\pm)}_{\mu\nu}=\pm\Sigma^{(\pm)}_{\mu\nu}$. The Euler-Lagrange variational equations for this functional were written down in [migdal2025SQYMflow]. These nonlinear differential equations were solved in that paper only for a special case of planar boundary loop $C$. In this paper, we generalize that solution for an arbitrary loop.
Self-dual area derivative. There is an internal $O(4)$ gauge invariance of our area element in addition to a physical $O(4)$ invariance of the problem in $\mathbb{R}^{4}$. We use that extra degree of freedom to provide the self-duality (SD or ASD) of the area derivative [migdal2025SQYMflow]. The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
$\frac{\delta W}{\delta\sigma_{\mu\nu}(\theta)}=$
$\frac{\delta}{\delta C^{\prime}_{\mu}(\theta-0)}\frac{\delta W}{\delta C^{\prime}_{\nu}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\mu}(\theta+0)}\frac{\delta W}{\delta C^{\prime}_{\nu}(\theta-0)}$ (2.3)
The loop equation for the Yang–Mills gradient flow was shown in [migdal2025SQYMflow] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
$\mathcal{L}(S)=0;$ (2.4)
$\mathcal{L}=\oint_{\theta}C^{\prime}_{\nu}(\theta)\left(\frac{\delta}{\delta C^{\prime}_{\mu}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\mu}(\theta-0)}\right)\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}$ (2.5)
The basic observation made in that paper was there is a SD solution similar to the Yang–Mills multi-instanton. The loop space diffusion equation (2.5) is identically satisfied for any functional $S[C]$ with SD or ASD area derivative
$\ast\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)}=\pm\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)};$ (2.6)
$e_{\alpha\mu\nu\lambda}\left(\frac{\delta}{\delta C^{\prime}_{\alpha}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\alpha}(\theta-0)}\right)\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)}\equiv 0;$ (2.7)
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [migdal2025SQYMflow] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents SD or ASD tensor. This would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
3 Geometric Definition of the Physical Area
The construction of the action $S_{\chi}$ in Sec. 4 depends on a specific choice of boundary conditions (BCs), which introduces a rotational ambiguity. The resulting functional is not $O(4)$ invariant. We must therefore provide a geometric definition of the physical area $S_{\mathrm{phys}}[C]$ that is fully invariant.
We define $S_{\mathrm{phys}}[C]$ as the solution of the minimax problem: first, we minimize the value of the Hodge-dual action $S[X]$ (2.1), by solving the Euler-Lagrange (E-L) equations for $S$ with specific boundary conditions $X=Q.C$ using a normalized matrix $Q$ compatible with the self-duality of the area derivative (see below). Then, we maximize this minimal value over the full group of allowed isometries of $Q$ (transformations that leave the E-L equations invariant).
Case 1: Simple (non-intersecting) loops
For a single, non-intersecting loop $C$, the E-L equations and local Dirichlet BCs $X=QC$ are global. The only relevant isometry that preserves the problem is a global $O(4)$ rotation, $R$. The rotation of $Q$ is equivalent to the opposite rotation of $C$ by the invariance of our area integral. The set of compatible solutions is thus $\{X[RC]\mid R\in O(4)\}$. The definition of the area becomes:
$S_{\mathrm{phys}}[C]:=\max_{R\in O(4)}S[X[RC]]$ (3.1)
This is precisely the definition that will be evaluated in Sec. 5, which yields the $O(4)$-invariant spectral law.
Case 2: Composite (self-intersecting) loops
For a loop that is a composite of two (or more) separated or self-intersecting sub-loops, $C=C_{1}\uplus C_{2}$, the problem changes. The E-L equations and the Dirichlet BCs are local. The problem of finding the minimal surface $X$ for the boundary $C$ therefore decouples into two independent minimal surface problems: $X_{1}$ on domain $\mathcal{D}_{1}$ bounded by $Q_{1}C_{1}$, and $X_{2}$ on domain $\mathcal{D}_{2}$ bounded by $Q_{2}C_{2}$. The total action for the solution is the sum $S[X]=S[X_{1}]+S[X_{2}]$.
Because the problem decouples, the isometry group of the total solution is larger. It is not the "diagonal" $O(4)$ group (which would force $R_{1}=R_{2}$), but the full product group $O(4)_{1}\times O(4)_{2}$. A transformation $(R_{1},R_{2})$ with $R_{1}\neq R_{2}$ is a valid isometry of the decoupled E-L equations.
To find the true maximum over "all compatible isometries" (per our definition), we must maximize over this larger group. This defines the area for a composite loop as $S_{\mathrm{phys}}[C_{1}\uplus C_{2}]=\max_{R_{1}\in O(4)}S[X_{1}[R_{1}C_{1}]]+\max_{R_{2}\in O(4)}S[X_{2}[R_{2}C_{2}]]=S_{\mathrm{phys}}[C_{1}]+S_{\mathrm{phys}}[C_{2}]$.
Finally, this maximized minimal projection area depends on the parametrization of the bounding curve, unlike the Wilson loop which is reparametrization invariant by definition. This last ambiguity is eliminated in the same way as in the Euclidean minimal surface: by minimizing the resulting functional over the group of orientation-preserving diffeomorphisms of the circle, $C(\theta)\to C(\phi(\theta))$ with $\phi^{\prime}>0$. Combining these steps, the full definition for a loop decomposable into sub-loops $C_{i}$ is:
$S[C]=\sum_{C_{i}\in C}$
$\min_{\phi\in\mathrm{Diff}^{+}(S^{1})}\max_{R\in O(4)}\min_{\begin{subarray}{c}X:\mathbb{D}_{i}\to\mathbb{R}^{4}\otimes\mathbb{R}^{4}\\ X|_{\partial\mathbb{D}_{i}}=QR(C_{i}\circ\phi)\end{subarray}}\int_{\mathbb{D}_{i}}\lVert dX^{A}\wedge dX^{A}\rVert$ (3.2)
4 A Self-Dual Holomorphic Ansatz
Holomorphic map.
We work in the target complex basis $z_{1}=x_{1}+\mathrm{i}x_{2}$, $z_{2}=x_{3}+\mathrm{i}x_{4}$. Let $U^{A}_{z_{a}}(z)$ be the holomorphic $(1,0)$ components of the embedding $X^{A}_{\mu}(z,\bar{z})=U^{A}_{\mu}(z)+\overline{U^{A}_{\mu}(z)}$. We also define $V^{A}_{z_{a}}:=\partial_{z}U^{A}_{z_{a}}$. As we shall see, the EL equations for the minimal Hodge-dual surface are identically satisfied by virtue of the Cauchy-Riemann equations for the holomorphic functions $U^{A}_{\mu}(z)$ combined with the SD property of the area element.
Lemma 4.1 (Normalization).
With the canonical SD $(1,1)$ form $B:=\mathrm{i}\big(dz_{1}\!\wedge d\bar{z}_{1}+dz_{2}\!\wedge d\bar{z}_{2}\big)$, the Euclidean Hodge star gives $\sqrt{\tfrac{1}{2}\,B_{\mu\nu}B_{\mu\nu}}=2\sqrt{2}$. Consequently, the area density of any SD surface $\Sigma=a(z)B$ is
$\sqrt{\tfrac{1}{2}\,\Sigma^{2}}\;$ = $=\;2\sqrt{2}\;a(z)$
$=\;\sqrt{2}\sum_{A}\big(|V^{A}_{z_{1}}|^{2}+|V^{A}_{z_{2}}|^{2}\big),$ (4.1)
, where the second line follows from $\Sigma_{z_{1}\bar{z}_{1}}=\Sigma_{z_{2}\bar{z}_{2}}=\mathrm{i}a(z)$ and the definition $\Sigma_{z_{a}\bar{z}_{b}}=\mathrm{i}\sum_{A}V^{A}_{z_{a}}\overline{V^{A}_{z_{b}}}$.
Proof.
In the real coframe $e^{1}=dx_{1},\dots,e^{4}=dx_{4}$, $\mathrm{i}dz_{1}\!\wedge d\bar{z}_{1}=2\,e^{1}\!\wedge e^{2}$ and $\mathrm{i}dz_{2}\!\wedge d\bar{z}_{2}=2\,e^{3}\!\wedge e^{4}$, so $B=2(e^{1}\!\wedge e^{2}+e^{3}\!\wedge e^{4})$. Then $\frac{1}{2}B^{2}=8$ and the identities follow. ∎
Two–form and Self-Duality.
The two-form is $\Sigma=\mathrm{d}U\wedge\mathrm{d}\overline{U}$. As $U$ is holomorphic, $\Sigma$ is of pure $(1,1)$ type. The self-duality condition $\ast\Sigma=\Sigma$ requires $\Sigma_{z_{1}\bar{z}_{1}}=\Sigma_{z_{2}\bar{z}_{2}}$. Using the component definition $\Sigma_{z_{a}\bar{z}_{b}}=\mathrm{i}\sum_{A}V^{A}_{z_{a}}\overline{V^{A}_{z_{b}}}$ from Lemma 4.1, this condition implies the explicit constraint:
$\sum_{A}\big(|V^{A}_{z_{1}}|^{2}\big)=\sum_{A}\big(|V^{A}_{z_{2}}|^{2}\big).$ (4.2)
The area density is given by Eq. (4.1).
An Explicit SD Ansatz.
We propose a minimal ansatz that satisfies the self-duality condition (4.2) using two independent holomorphic functions $f_{1}(z)$ and $f_{2}(z)$. Let the only non-zero components of $U$ be:
$U^{4}_{z_{1}}=f_{1}(z),\hskip 18.49988ptU^{3}_{z_{2}}=f_{1}(z),$
$U^{2}_{z_{1}}=f_{2}(z),\hskip 18.49988ptU^{1}_{z_{2}}=f_{2}(z).$ (4.3)
This choice manifestly satisfies (4.2), as $\sum_{A}|V^{A}_{z_{1}}|^{2}=|f_{1}^{\prime}|^{2}+|f_{2}^{\prime}|^{2}$ and $\sum_{A}|V^{A}_{z_{2}}|^{2}=|f_{2}^{\prime}|^{2}+|f_{1}^{\prime}|^{2}$.
Boundary Conditions and Explicit Projector.
The action functional (4.1) for this ansatz becomes a Dirichlet functional
$S_{\chi}=2\sqrt{2}\int_{\mathcal{D}}\big(|f_{1}^{\prime}|^{2}+|f_{2}^{\prime}|^{2}\big)\mathrm{d}^{2}z.$ (4.4)
The EL equations of this functional are satisfied by Cauchy-Riemann equations for $f_{1},f_{2}$, so that this functional reduces to the boundary data. By the Dirichlet identity, $S_{\chi}=2\sqrt{2}\sum_{n\geq 1}n\big(|\hat{f}_{1,n}|^{2}+|\hat{f}_{2,n}|^{2}\big)$. We now fix the boundary conditions by relating $f_{1},f_{2}$ to the boundary loop $C_{\mu}$ in a rotated basis:
$\Re f_{1}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{1}(\theta)+C_{2}(\theta)\big),$
$\Re f_{2}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{3}(\theta)+C_{4}(\theta)\big).$ (4.5)
Substituting the Fourier modes $\hat{f}_{1,n}=\frac{1}{\sqrt{2}}(\hat{C}_{n,1}+\hat{C}_{n,2})$ and $\hat{f}_{2,n}=\frac{1}{\sqrt{2}}(\hat{C}_{n,3}+\hat{C}_{n,4})$ into (4.4):
$S_{\chi}$ $=2\sqrt{2}\sum_{n\geq 1}n\left(\Big|\frac{\hat{C}_{n,1}+\hat{C}_{n,2}}{\sqrt{2}}\Big|^{2}+\Big|\frac{\hat{C}_{n,3}+\hat{C}_{n,4}}{\sqrt{2}}\Big|^{2}\right)$
$=\sqrt{2}\sum_{n\geq 1}n\left(|\hat{C}_{n,1}+\hat{C}_{n,2}|^{2}+|\hat{C}_{n,3}+\hat{C}_{n,4}|^{2}\right).$ (4.6)
Introducing the Douglas-Gram matrix $G_{\mu\nu}:=\sum_{n\geq 1}n\,\hat{C}_{n,\mu}\,\overline{\hat{C}_{n,\nu}}$, this action is a quadratic form $S_{\chi}=2\sqrt{2}\,\operatorname{tr}(PG)$, where $P$ is the explicit rank-2 projector $P=\frac{1}{2}\left(\begin{smallmatrix}1&1&0&0\\ 1&1&0&0\\ 0&0&1&1\\ 0&0&1&1\end{smallmatrix}\right)$.
Lemma 4.2 (Projector Property).
The matrix $P$ defined by (4) is a rank-2 orthogonal projector ($P\in\operatorname{Gr}(2,4)$). This is verified by direct computation: $P$ is symmetric, $P^{2}=P$, and $\operatorname{tr}P=2$, so $\operatorname{rank}(P)=2$.
Regular loops.
Throughout we assume the boundary loop $C:S^{1}\to\mathbb{R}^{4}$ is regular in the sense that the spectral expansion converges in the Douglas–Gram matrix, so that the Dirichlet identity leading to (4.6) is valid, and that the Fourier/Hilbert-transform manipulations used in this paper are justified. This corresponds to $C\in H^{\tfrac{1}{2}}(S^{1},\mathbb{R}^{4})$
5 Invariant Area and the Spectral Law
Rotationally invariant area.
The true, physical $O(4)$-invariant area $S_{\mathrm{phys}}$ was defined as the maximum area that can be generated by our functional form over all possible orientations $R$ of the loop $C$. For this holomorphic solution
$S_{\mathrm{phys}}[\mathcal{C}]$ $\;:=\;\max_{R\in O(4)}S_{\chi}[RC;P]$
$\;=\;\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}\!\left(P\,(RGR^{\top})\right).$ (5.1)
Let $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$ be the eigenvalues of $G$. The maximum in (5.1) is evaluated by first applying the cyclic property of the trace, which gives $\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}((R^{T}PR)G)$. We can define a new projector $P^{\prime}=R^{T}PR$. Since $P$ is a fixed rank-2 projector, $P^{\prime}$ is also a rank-2 projector. The group $O(4)$ acts transitively on the space of all 2-planes, $\operatorname{Gr}(2,4)$. This means that by maximizing over all $R\in O(4)$, the projector $P^{\prime}$ sweeps through the entire set of all possible rank-2 projectors. Therefore, our maximization is mathematically identical to $2\sqrt{2}\max_{P^{\prime}\in\operatorname{Gr}(2,4)}\operatorname{tr}(P^{\prime}G)$. By the Ky Fan theorem [Horn:2012:MatrixAnalysis:2nd], this maximum is precisely the sum of the two largest eigenvalues, $\lambda_{3}+\lambda_{4}$.
Outcome and the spectral law. Crucially, the result of this maximization is $S_{\mathrm{phys}}[\mathcal{C}]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})$. This value is a universal spectral invariant, independent of the specific rank-2 projector $P$ we started with.
$S_{\mathrm{phys}}[\mathcal{C}]\;=\;2\sqrt{2}\,(\lambda_{3}+\lambda_{4}).$ (5.2)
This calculation confirms that our physical definition of the invariant area (5.1) leads directly to the $O(4)$-invariant spectral law. For planar loops ($\operatorname{rank}G=2$), $\lambda_{1}=\lambda_{2}=0$, and this correctly specializes to $S_{\mathrm{phys}}[\mathcal{C}]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})=2\sqrt{2}\,\mathcal{D}[C]$, reproducing the claim of Theorem 1 of [migdal2025SQYMflow]. The optimal projector $P_{\ast}=v_{3}v_{3}^{\top}+v_{4}v_{4}^{\top}$ (where $v_{i}$ are eigenvectors of $G$) is the one that achieves this maximum.
5.1 Projector Alignment
The physical area $S_{phys}$ is realized by maximizing the base functional $S_{\chi}$ over the global $\mathrm{O}(4)$ rotation group $\mathbf{R}$: $S_{phys}=\max_{\mathbf{R}\in\mathrm{O}(4)}2\sqrt{2}\operatorname{tr}(\mathbf{P}(\mathbf{RGR}^{\top}))$. The initial fixed projector $\mathbf{P}$ is defined by the orthonormal basis vectors $\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ associated with the Hodge parity $\chi=\pm 1$:
$\mathbf{u}_{1}$ $=\frac{1}{\sqrt{2}}(\mathbf{e}_{1}+\mathbf{e}_{2})$ (5.3)
$\mathbf{u}_{2}$ $=\frac{1}{\sqrt{2}}(\mathbf{e}_{3}+\chi\mathbf{e}_{4})$ (5.4)
The optimal rotation $\mathbf{R}_{max}=\mathbf{R}_{2}\mathbf{R}_{1}$ maps this initial SD basis onto the optimal basis $\{\mathbf{v}_{3},\mathbf{v}_{4}\}$ corresponding to the largest eigenvalues of $\mathbf{G}$, $\lambda_{3}$ and $\lambda_{4}$.
5.1.1 1. Projection and Orthogonalization (the combined operation $\mathbf{R}_{1}$
The first operation $\mathbf{R}_{1}$ starts with alignement of the initial plane $\operatorname{span}(\mathbf{u}_{1},\mathbf{u}_{2})$ with the optimal eigenspace $\mathcal{V}_{34}=\operatorname{span}(\mathbf{v}_{3},\mathbf{v}_{4})$. This is achieved by projecting $\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ orthogonal to the null space spanned by the two lowest eigenvectors $\{\mathbf{v}_{1},\mathbf{v}_{2}\}$. Let $\mathbf{P}_{12}=\mathbf{v}_{1}\mathbf{v}_{1}^{\top}+\mathbf{v}_{2}\mathbf{v}_{2}^{\top}$ be the projector onto this null space. The intermediate vectors $\mathbf{u}^{\prime}_{i}$ are given by: $\mathbf{u}^{\prime}_{i}=\operatorname{Normalize}\left[\mathbf{u}_{i}-\mathbf{P}_{12}\cdot\mathbf{u}_{i}\right]$
The second part of this operation $\mathbf{R}_{1}$ orthogonalizes these two vectors: $\mathbf{u}^{\prime\prime}_{2}=\operatorname{Normalize}\left[\mathbf{u}_{2}-(\mathbf{u}_{2}\cdot\mathbf{u}_{1})\mathbf{u}_{1}\right]$ After these two transformations, the basis vectors $\mathbf{u}_{1},\mathbf{u}_{2}$ become orthogonal basis in the Eigenspace ($\mathbf{R}_{2}$. Therefore, this combined transformation is an $O(4)$ rotation.
5.1.2 2. Alignment within the Eigenspace ($\mathbf{R}_{2}$)
The second rotation $\mathbf{R}_{2}$ performs an $\mathrm{O}(2)$ transformation within the optimal plane $\mathcal{V}_{34}$ to fix the alignment of the resulting basis. This step aligns the first intermediate vector $\mathbf{u}^{\prime}_{1}$ with the highest eigenvector $\mathbf{v}_{3}$: $\mathbf{R}_{2}:\quad\text{An }\mathrm{O}(2)\text{ rotation in }\mathcal{V}_{34}\text{ such that }\mathbf{R}_{2}\mathbf{u}^{\prime}_{1}\propto\mathbf{v}_{3}$ The combined rotation $\mathbf{R}_{max}$ defines the required boundary condition $\mathbf{X}=\mathbf{Q}\cdot\mathbf{R}_{max}\mathbf{C}$ for the minimal surface, which yields the spectral area $S_{phys}$.
6 Self-Duality of the Area Derivative
Theorem 6.1 (Self-Duality of the Physical Action).
The physical action $S_{\mathrm{phys}}[\mathcal{C}]$ defined in (5.2) has a self-dual (SD) area derivative:
$\ast\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}=\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}.$ (6.1)
Proof.
The physical action $S_{\mathrm{phys}}[\mathcal{C}]$ is the envelope (the maximum) of a family of functionals $S_{R}[C]:=S_{\chi}[RC;P]$, indexed by the rotation $R\in O(4)$, where $S_{\chi}[C;P]$ is the base functional from Sec. 4. The area derivative operator $\mathcal{D}:=\frac{\delta}{\delta\sigma}$ and the Hodge star $\ast$ are both $O(4)$-covariant. This means that if we can prove self-duality for any single member of this family (e.g., the $R=I$ case, $S_{\chi}[C;P]$), then self-duality holds for all members. The derivative of the envelope $S_{\mathrm{phys}}$ at a loop $C$ is, by definition, the derivative of the specific functional $S_{R_{\text{max}}}[C]$ that achieves the maximum at that point. Since all functionals in the family will be shown to have an SD derivative, $S_{\mathrm{phys}}$ must also have an SD derivative. Therefore, the proof reduces to verifying the self-duality for the base functional $S_{\chi}[C;P]$ (for $R=I$), which we do now.
We begin with the generic formula from [migdal2025SQYMflow] for the first variation of the Hodge-dual area
$\delta S=2\int d\theta\delta X^{A}_{\mu}(\theta)\partial_{\theta}X^{A}_{\nu}(\theta)T_{\mu\nu}$ (6.2)
where $T_{\mu\nu}$ is the normalized area element $\Sigma_{\mu\nu}$. [^1] This is generic variation of our functional in the space of all functions $X^{A}_{\mu}(u,v)$, evaluated on our holomorphic solution after taking a variation. Had we computed the variation on the class of holomorphic functions by reducing the Dirichlet functional to the boundary, we would have found a symmetric quadratic form with a vanishing (and trivial) area derivative. Thus, variations orthogonal to the solution manifold are essential for the non-trivial area derivative. With our holomorphic solution, using the constant $B_{\mu\nu}$ tensor and the Dirichlet functional with the Hilbert transform, we find
$\delta S\propto\iint d\theta_{1}d\theta_{2}H(\theta_{1}-\theta_{2})Q(C(\theta_{1}),\delta C^{\prime}(\theta_{2}));$ (6.3)
$Q(A,B)=A_{2}B_{1}-A_{1}B_{2}+A_{3}B_{4}-A_{4}B_{3}$ (6.4)
Now, the first and second variations are
$\frac{\delta S}{\delta C^{\prime}(\theta_{2})}\propto\int d\theta_{1}H(\theta_{1}-\theta_{2})Q(C(\theta_{1}),I_{4});$ (6.5)
$\left.\frac{\delta^{2}S}{\delta C^{\prime}_{\mu}(\theta^{\prime})\delta C^{\prime}_{\nu}(\theta)}\right|_{\theta-0}^{\theta+0}\propto\int_{0}^{\theta}d\theta^{\prime}H(\theta^{\prime}-\theta)\hat{Q}_{\mu\nu};$ (6.6)
where $\hat{Q}_{\mu\nu}$ is the constant, antisymmetric tensor $\hat{Q}=\text{diag}(i\sigma_{2},i\sigma_{2})$. In the second derivative, by the chain rule, we use the standard Volterra derivative [Gelfand:1963:CalculusVariations] of the loop with respect to its tangent, $\delta C(x)/\delta C^{\prime}(y)=\Theta(x-y)$. This tensor $\hat{Q}$ is manifestly self-dual (as $\ast\hat{Q}_{12}=\hat{Q}_{34}$ and $\ast\hat{Q}_{34}=\hat{Q}_{12}$), which proves that the area derivative for $S_{\chi}[C;P]$ is self-dual. By the covariance argument above, this property extends to $S_{\mathrm{phys}}$.
Note that the numerical factor $\int_{0}^{\theta}d\theta^{\prime}H(\theta^{\prime}-\theta)\propto\sum_{n>0}1/n$ logarithmically diverges. However, the SD of the area derivative, valid for the tensor factor $Q_{\mu\nu}$, leads to the cancelation of the whole series in the loop space diffusion equation (2.5) (as argued in Sec. 2). Note that this cancelation is valid irrespective of the cutoff at large $n$ in Fourier series, so that if we fix this cutoff and substitute the area derivative into the loop equation (2.5), we get exactly zero. Thus, we can still claim the finite solution of the loop equation , as the minimal area is finite, and its contribution to the loop space diffusion equation is zero, making it a fixed point. ∎
7 The Hodge parity flip
How does our solution change for the negative Hodge parity $\chi=-1$ (ASD instead of SD)? There are many equivalent forms of the ASD solutions, all equivalent after the isometry transformation $O(4)$ rotation of the matrix $Q\Rightarrow QR$ in the Dirichlet BC $X=QC$. The conical transformation is a reflection $X^{A}_{4}\Rightarrow-X^{A}_{4}$ which in complex basis corresponds to $z_{2}\leftrightarrow\bar{z}_{2}$ in our holomorphic ansatz (4.3). After reflection, the only nonzero components will be
$U^{4}_{z_{1}}=f_{1}(z),\hskip 18.49988ptU^{3}_{\bar{z}_{2}}=f_{1}(z),$
$U^{2}_{z_{1}}=f_{2}(z),\hskip 18.49988ptU^{1}_{\bar{z}_{2}}=f_{2}(z).$ (7.1)
As a result, the $(2,0)$ and $(0,2)$ components of the form $\Sigma$ will still be zero, and the $(0,1),(1,0)$ components change. Now the $\Sigma$ form will become $\Sigma=a(z)\tilde{B}$, where $\tilde{B}~e_{1}\wedge e_{2}-e_{4}\wedge e_{4}$ which has negative Hodge duality. The energy density $a(z)$ stays the same, and so does the area functional. The BC (4.5) change in an obvious way
$\Re f_{1}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{1}(\theta)+C_{2}(\theta)\big),$
$\Re f_{2}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{3}(\theta)-C_{4}(\theta)\big).$ (7.2)
which leaves the projection operator $P$ unchanged.
The area derivative in (6.4) will change in an obvious way, following the change of the area element $B\Rightarrow\tilde{B}$
$\delta S\propto\iint d\theta_{1}d\theta_{2}H(\theta_{1}-\theta_{2})\tilde{Q}(C(\theta_{1}),\delta C^{\prime}(\theta_{2}));$ (7.3)
$\tilde{Q}(A,B)=A_{2}B_{1}-A_{1}B_{2}-A_{3}B_{4}+A_{4}B_{3}$ (7.4)
As expected, the area does not depend on the Hodge parity, but the area derivative switches from SD to ASD under Hodge parity flip. The loop equation will still hold, the same way as for the SD solution.
8 Reparametrization and additivity
Parametric invariance. As stated below (5.1) and in Sec. 3, the geometric area is defined by maximization over allowed isometries; we additionally take the infimum over orientation–preserving reparametrizations to remove any dependence on the boundary parameter. We therefore adopt the parametrization–invariant form
$S_{\mathrm{phys}}[C]$
$:=\min_{\phi\in\mathrm{Diff}^{+}(S^{1})}\;\max_{R\in O(4)}\,2\sqrt{2}\;\operatorname{tr}\!\Big(P\,R\,G[C\!\circ\!\phi]\,R^{\top}\Big).$ (8.1)
By construction, the choice of parameter is without loss of generality: one may first reparametrize to a convenient standard $0,2\pi]$ and only then perform the $O(4)$ maximization and the outer minimum over $\phi$.
Canonical split at a self–intersection. Let $C(\theta_{1})=C(\theta_{2})$ be a transversal self–intersection and define the two oriented subloops $C_{1}=C[\theta_{1},\theta_{2}]$ and $C_{2}=C[\theta_{2},\theta_{1}\!+\!2\pi]$ as in Sec. 3 (Case 2). Using parametric invariance, choose $\phi_{1},\phi_{2}\in\mathrm{Diff}^{+}(S^{1})$ with $\phi_{1}(S^{1})=[\theta_{1},\theta_{2}]$ and $\phi_{2}(S^{1})=[\theta_{2},\theta_{1}\!+\!2\pi]$ and set $\widetilde{C}_{j}(\tau)=C(\phi_{j}(\tau))$ so that each boundary component is on $0,2\pi]$. For a disconnected boundary the Douglas–Gram series splits,
$G[\widetilde{C}_{1}\!\uplus\!\widetilde{C}_{2}]$ $=G[\widetilde{C}_{1}]+G[\widetilde{C}_{2}],$ (8.2)
and, as explained in Sec. 3 (Case 2), the E–L problems decouple and the isometry group enlarges to the product $O(4)_{1}\times O(4)_{2}$.
Additivity of $S_{\mathrm{phys}}$. With (8.2), the double extremization factorizes:
$S_{\mathrm{phys}}[C_{1}\!\uplus\!C_{2}]:=2\sqrt{2}\,$
$\min_{\phi_{1},\phi_{2}}\;\max_{\begin{subarray}{c}R_{1}\in O(4)\\ R_{2}\in O(4)\end{subarray}}\Big\{\operatorname{tr}\!\big(P\,R_{1}G[C_{1}\!\circ\!\phi_{1}]R_{1}^{\top}\big)$
+ $\hskip 73.07445pt+\,2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{2}G[C_{2}\!\circ\!\phi_{2}]R_{2}^{\top}\big)\Big\}$ (8.3)
$=\min_{\phi_{1}}\max_{R_{1}}2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{1}G[C_{1}\!\circ\!\phi_{1}]R_{1}^{\top}\big)$
$+\;\min_{\phi_{2}}\max_{R_{2}}2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{2}G[C_{2}\!\circ\!\phi_{2}]R_{2}^{\top}\big)$
$=\;S_{\mathrm{phys}}[C_{1}]+S_{\mathrm{phys}}[C_{2}].$ (8.4)
The first equality is the definition consistent with Sec. 3 (Case 2); the second uses the product structure of the objective (sum of two independent functionals), which makes the $\max$ over $R_{1},R_{2}$ and the $\min$ over $\phi_{1},\phi_{2}$ separate.
Compatibility with the MM kernel. The right–hand side of the MM equation [MMEq79, MM1981NPB] is supported on self–intersections via the bilocal kernel $\int\!\!\int d\theta_{1}d\theta_{2}\,C^{\prime}(\theta_{1})\!\cdot\!C^{\prime}(\theta_{2})\,\delta(C(\theta_{1})-C(\theta_{2}))\,\cdots$. Under $\theta_{a}=\phi_{a}(\tau_{a})$ (with $\phi_{a}^{\prime}\!>\!0$),
$d\theta_{a}\,C^{\prime}(\theta_{a})$ $=d\tau_{a}\,\dot{C}(\tau_{a}),$
$\delta\!\big(C(\theta_{1})-C(\theta_{2})\big)$ $=\delta\!\big(C(\phi_{1}(\tau_{1}))-C(\phi_{2}(\tau_{2}))\big),$ (8.5)
So, the kernel is parameterization-invariant. Therefore, the canonical standardization to $0,2\pi]$ used above is legitimate and (8.4) guaranties that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\uplus\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.5) and [Mig83, migdal2025SQYMflow]).
Let us stress that this factorization takes place independently of the splitting points $\theta_{1},\theta_{2}$. Thus, for every $\theta_{1},\theta_{2}$, the product of two factors $\exp\left(-\kappa S[C_{1}\right)$ and $\exp\left(-\kappa S[C_{2}]\right)$ exactly equals the factor $\exp\left(-\kappa S[C]\right)$ on the left side of the MM equation; thus, it can be taken out of the integral, after which it cancels on both sides of the MM equation.
9 Confinement theorem
Theorem 9.1 (Confinement area law).
Let $G$ be the Douglas–Gram matrix of $C$ with eigenvalues $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$. as we have proven in the previous section, the Hodge dual area and the Euclidean minimal area satisfy
$A_{\mathrm{geom}}[C]:=\operatorname{tr}G\;=\;\sum_{i=1}^{4}\lambda_{i},\qquad S_{\chi}[C]\;=\;2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$ (9.1)
$\sigma_{\mathrm{eff}}[C]:=\frac{S_{\chi}[C]}{A_{\mathrm{geom}}[C]}\;=\;2\sqrt{2}\,\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}.$ (9.2)
Moreover,
[…]
with equality on the right for planar spectra ($\lambda_{1}=\lambda_{2}=0$),and on the left for isotropic spectra ($\lambda_{1}=\cdots=\lambda_{4}$).
The quantity $\sigma_{\mathrm{eff}}[C]$ (hence $S_{\chi}[C]$) is O(4)O(4)–invariant and chirality–even: if $C\mapsto RC$ with $R\in O(4)$, then $G\mapsto RGR^{\top}$ leaves $\{\lambda_{i}\}$, and thus the theorem, unchanged.
Remark 9.1 (Minimization over reparametrization).
Both functionals, Douglas’s $\sum\lambda_{i}$ and our $\lambda_{3}+\lambda_{4}$, have yet to be minimized over the reparametrization of the bounding loop $C(\theta)\Rightarrow C(f(\theta)),f^{\prime}>0$. As the minimization of just the two highest eigenvalues, compared to that of the trace, is attained at different parameterizations, we can only state that $\sigma_{eff}=2\sqrt{2}\frac{\min(\lambda_{3}+\lambda_{4})}{\min(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})}$. However, the lower inequality still holds $\sigma_{eff}\geq\sqrt{2}$, as for every parametrization $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}\leq 2(\lambda_{3}+\lambda_{4})$.Therefore, the minimum in the denominator is less than or equal to twice that in the numerator, proving the lower bound quoted in confinement theorem.
10 Master Field and the Meaning of the Minimal Surface in Planar QCD
Planar QCD represents the large-$N_{c}$ limit of a Quantum Field Theory defined by gauge fields, summing over all planar graphs to describe fluctuations around a (yet unknown) vacuum state. However, when formulated in terms of the Wilson loop observable $\Psi[C]=\langle\frac{1}{N}\mathrm{Tr}\mathcal{P}\exp\left(\oint_{C}Adx\right)\rangle$, the theory undergoes a fundamental shift: it ceases to be a quantum theory and becomes a classical field theory in loop space.
In the large $N_{c}$ limit, due to factorization $\langle W(C_{1})W(C_{2})\rangle=\langle W(C_{1})\rangle\langle W(C_{2})\rangle+O(1/N^{2})$, the functional $\Psi[C]$ becomes the Master Field, satisfying a deterministic nonlinear equation— the MM loop equation.
It is crucial to understand that in this framework, we are not “quantizing a string theory” in the traditional sense. Rather, we are solving a classical field equation in loop space. The introduction of the Hodge-dual minimal surface in the auxiliary space $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ is a method to construct this solution.
This approach mirrors the logic of the Eguchi-Kawai (EK) reduction [Das:1987dg]. In the EK model, quenched “momenta” or twisted boundary conditions are not introduced as inherent physical degrees of freedom of the vacuum, but as mathematical instruments designed to satisfy the loop equations of the planar theory on a single lattice site. Similarly, in our construction, the only physical requirement is the satisfaction of the MM equation; the string analogy applies strictly insofar as it serves this goal.
There is no claim of a direct physical connection between the microscopic QCD gluon variables and the embedding coordinates $X_{\mu}^{A}$ in the 16-dimensional auxiliary space. Unlike the AdS/CFT correspondence, we do not postulate a full quantum equivalence between planar QCD and a fundamental string theory in a higher-dimensional bulk. Instead, we construct the specific additive zero mode of the loop equation, in the same spirit as one solves for the spontaneous magnetization in a ferromagnet: identifying the non-perturbative classical configuration that creates the confining background.
Still, the technology of String Theory, if not its complete physical contents, can be quite useful in elaborating this solution of planar QCD.
11 Topological String with Nonlinear Endpoint Dynamics
Our area functional is equivalent to the extremal value of the string theory action
$S_{0}[f_{i}]$ $=\int_{\mathbb{D}}d^{2}z\,(\nabla f_{i})^{2}+\int_{\mathbb{D}}d^{2}z\,u_{i}\,\partial_{\bar{z}}f_{i},$ (11.1)
where $f_{1,2}(z)$ are the basis functions introduced in Section 4. On the boundary, we have linear boundary conditions of the form:
$S_{\partial}[f_{i},\rho]$ $=\oint d\theta\,\rho_{i}(\theta)\left[\Re f_{i}\left(e^{i\theta}\right)-Q_{i}(C(\theta))\right],$ (11.2)
with $Q_{1}(C)=(C_{1}+C_{2})/\sqrt{2}$ and $Q_{2}(C)=(C_{3}+\chi C_{4})/\sqrt{2}$, where $\chi=\pm 1$ is the Hodge parity. Integrating out the auxiliary fields $f_{i}$ reproduces the Douglas-type boundary action $\mathcal{D}[QC]=\mathrm{Tr}(PG)$.
This system describes a topological field theory: the action reduces entirely to boundary values as a consequence of the imposed Cauchy-Riemann equations. Physically, this action represents the net area of two conformal maps, $f_{1}(\mathbb{D})$ and $f_{2}(\mathbb{D})$, with no remaining bulk degrees of freedom. In string theory language, there are no vibrations of the string body; there are only the nonlinear dynamics of its endpoints determined by the loop shape. In this sense, it is analogous to 2D quantum gravity, where the Einstein-Hilbert action reduces to the Euler characteristic, independent of the local geometry in the bulk.
The self-duality of the area derivative stems from local Hodge duality and holds quantum-mechanically. The quadratic nature of the action ensures that the quantum Schwinger–Dyson equations match the classical Euler-Lagrange relations. Thus, the expectation value $\langle\delta S/\delta\sigma\rangle$ is self-dual, and the quantum string obeys the classical loop equations.
Integrating out $f_{i}$ yields a Gaussian integral:
$Z[C;P]$ $\propto[\det{}^{\prime}(K_{P})]^{-1/2}$
$\times\exp\left(-2\sqrt{2}\,\kappa\mathrm{Tr}(PG)\right),$ (11.3)
where $K_{P}=P\,H\,\partial_{\theta}P$ has a universal spectrum. Hence $\det{}^{\prime}(K_{P})$ is a constant, and the effective action is classical: $\Gamma[C;P]=2\sqrt{2}\,\mathrm{Tr}(PG)+\text{const}$.
According to our min-max-min principle, which is required to preserve the factorization of the product of dressing factors $\exp\left(-\kappa S[C]\right)$ on the right side of the MM equation, we do not integrate over the “moduli space” (the rotation matrix $R$). Instead, we select the optimal rotation that brings the exponent to its maximal value:
$Z[C]$ $\propto\exp\left(-2\sqrt{2}\,\kappa\max_{R\in SO(4)}\mathrm{Tr}(R^{T}PRG)\right)$
$=\exp\left(-2\sqrt{2}\,\kappa(\lambda_{3}+\lambda_{4})\right).$ (11.4)
This corresponds to setting an optimal isometry of the boundary condition rather than averaging over these isometries. In terms of the partition function $Z=\exp\left(-\kappa S[RC]\right)$, our prescription selects the minimal value, rather than performing a saddle point integration around the maximal value of this factor.
Let us stress that we are solving the classical loop equations of large-$N_{c}$ QCD rather than quantizing a string theory. The internal degrees of freedom found in non-critical string theory (such as Polyakov’s Liouville mode) do not apply here, as our theory is based on the classical dynamics of Wilson loop trace functionals in planar QCD. We utilize string theory as a mathematical tool, rather than as a physical object describing the condensation of electric flux in the QCD vacuum.
In our picture, the fluctuations of this electric flux are described by planar graphs for gluons in another factor $W_{pert}[C]$. The confining factor $Z[C]$ only reflects the mean properties of the vacuum, frozen in the large $N_{c}$ limit. The optimal rotation $R_{*}$ described in section 5.1, aligning the SD/ASD projector with the highest plane of the Gram basis determines the boundary conditions for the holomorphic string:
$\Re f_{i}\left(e^{i\theta}\right)=Q_{i}(R_{*}C(\theta)).$ (11.5)
For the ASD string, the reflected coordinate $C_{4}$ is used in $Q_{i}(RC)$.
12 String Scale $\kappa$ and Gluon Condensate
The string mass scale $\kappa$ is not fixed by the MM equation itself. As we have seen, the confining factor $\exp\left(-\kappa S[C]\right)$ passes through the equation without correlation with the perturbative solution $W_{pert}[C]$. This factor reflects the non-perturbative properties of the QCD vacuum rather than the perturbative fluctuations of the gluon field.
The fundamental non-perturbative property of the QCD vacuum is the gluon condensate [Shifman:1978bx]. We relate the string tension $\kappa$ to this condensate by considering a circular loop $C$ of small radius $r$. According to our solution, $S[C]=2\sqrt{2}\pi r^{2}$. We compute the second area derivative at two opposite points $x_{1}=C(0)$ and $x_{2}=C(\pi)$ on the circle.
At small distances $|x_{12}|=2r$, the Operator Product Expansion (OPE) in asymptotically free QCD is dominated by the leading conformal term in $W_{pert}[C]$:
$\frac{\delta^{2}W_{pert}[C]}{\delta\sigma_{\mu\nu}(1)\delta\sigma_{\lambda\rho}(2)}$ $\propto\langle\mathrm{Tr}F_{\mu\nu}(x_{1})U_{0,\pi}F_{\lambda\rho}(x_{2})U_{\pi,2\pi}\rangle$
$\to\frac{\alpha_{eff}}{4\pi}\frac{I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}}{|x_{12}|^{4}}W_{pert}[C],$ (12.1)
where $I_{\mu\lambda}=\delta_{\mu\lambda}-2\hat{x}_{\mu}\hat{x}_{\lambda}$ with $\hat{x}=x_{12}/|x_{12}|$. We project this perturbative singularity out by contracting with the tensor $\Pi_{\mu\nu\lambda\rho}=\tfrac{1}{2}(\delta_{\mu\lambda}\delta_{\nu\rho}-\delta_{\nu\lambda}\delta_{\mu\rho})$. Simple tensor algebra confirms that the leading term vanishes: $\mathrm{Tr}(\Pi I\otimes I)=0$.
The remaining term is the gluon condensate contribution. This term involves the renormalization factor $Z^{2}$ associated with the two cusps created at the points of differentiation:
$\Pi_{\mu\nu\lambda\rho}\mathrm{Tr}F_{\mu\nu}(1)U_{1,2}F_{\lambda\rho}(2)U_{2,1}\to$
$-\langle(G^{a}_{\mu\nu})^{2}\rangle W_{pert}[0]Z^{2}.$ (12.2)
This constant contribution must be balanced by the second area derivative of our Hodge-dual factor in the ansatz $W=W_{pert}e^{-\kappa S}$. Using the fact that the first derivative $\delta S/\delta\sigma$ vanishes for the circle due to symmetry, we find:
$W_{pert}[C]\frac{\delta^{2}\exp\left(-\kappa S[C]\right)}{\delta\sigma_{\mu\nu}(1)\delta\sigma_{\lambda\rho}(2)}\to$
$-\kappa W_{pert}[0]\Pi_{\mu\nu\lambda\rho}\frac{\delta^{2}S[C]}{\delta\sigma_{\mu\nu}(0)\delta\sigma_{\lambda\rho}(\pi)}.$ (12.3)
Crucially, there is no $Z$ factor in this geometric term; it depends only on the renormalization of the perturbative loop $W_{pert}[0]$ which appears as a common prefactor.
The second area derivative is computed as the second variation of $S=2\sqrt{2}(\lambda_{3}+\lambda_{4})$:
$\frac{\delta^{2}S[C]}{\delta\sigma^{2}}\propto\mathrm{Tr}(\Pi B\otimes B)\sum_{i<j}\frac{\left(\sum_{n>1}\frac{1}{n}\right)^{2}}{\lambda_{j}-\lambda_{i}}.$ (12.4)
The harmonic sum $\sum_{n>1}1/n\to\log n_{max}$ arises from the singular Hilbert kernels in the area derivatives. The result is:
$\frac{\delta^{2}S[C]}{\delta\sigma^{2}}=\frac{a(\log n_{max})^{2}}{r^{2}}.$ (12.5)
The sign matches the gluon condensate term, while the magnitude depends on the radius.
Remark 12.1.
The divergence scaling as $1/r^{2}$ arises from the second area derivative as a direct consequence of the scaling property of our Hodge minimal area. This is a nonlinear, nonlocal, homogeneous functional of degree two. Every area derivative involves two functional derivatives with respect to $\dot{C}$, lowering the degree by two. Thus, the second area derivative has degree minus two, inevitably scaling as $1/r^{2}$. The logarithmic divergence comes from the high-frequency spectral sum.
The proper physical procedure is to match these terms at the crossover scale. In the matching region, we equate the renormalized OPE result with the geometric derivative:
$Z^{2}\langle(G^{a}_{\mu\nu})^{2}\rangle=\frac{\kappa a(\log n_{max})^{2}}{r^{2}}\exp\left(-2\sqrt{2}\pi r^{2}\kappa\right).$ (12.6)
where the constant $a\sim 1$ is calculable from our solution, and renormalization constant $Z$ for the area derivative operator applied to $W_{pert}[C]$ is calculable by the methods of [KorchemskyRadyushkin1987]. Presumably, the renormalization factor $Z^{2}$ on the LHS compensates the UV divergence $(\log n_{max})^{2}$ on the RHS, rendering the relation finite and renormalization-group invariant.
The function $f(z)=ze^{-z}$ (where $z=2\sqrt{2}\pi r^{2}\kappa$) reaches its maximum at $z=1$. This implies that no solutions exist outside this parameter region, suggesting a critical point. We identify this maximum as the natural matching point:
$r_{c}^{2}=\frac{1}{2\sqrt{2}\pi\kappa};$ (12.7)
$\kappa^{2}=\frac{e}{2\sqrt{2}a\pi}\left(\frac{Z}{\log n_{max}}\right)^{2}\left\langle(G^{a}_{\mu\nu})^{2}\right\rangle.$ (12.8)
The factors of $N_{c}$ cancel, consistent with the planar limit. This matching fixes the string tension $\kappa$ in terms of the gluon condensate $\langle G^{2}\rangle$.
Our association of $\kappa$ with the gluon condensate supports the physical picture of condensed electric flux originally conjectured by Nambu [Nambu:1974zg]. One may view the term $\kappa S[C]$ as the action cost of a field configuration with energy density $\langle G^{2}\rangle\sim\Lambda_{QCD}^{4}$ spread over a “pancake” surface of thickness $\Lambda_{QCD}^{-2}$. This estimate interprets the dressing factor $\exp\left(-\kappa S[C]\right)$ as a statistical probability $\exp\left(-\beta\delta E\right)$ for a large quark loop propagating through the QCD vacuum (the dual Meissner effect).
13 Conclusion
We have presented an exact solution for the quaternionic Hodge–dual minimal surface, a new class of minimal surfaces in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ that is exactly solvable by reduction to free conformal maps. The minimized area is the non-trivial spectral invariant $S_{\min}(C)=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})$, where $\{\lambda_{i}\}$ are the eigenvalues of the Douglas–Gram matrix $G$. This result is mathematically robust: the optimal surface is determined by the two dominant Douglas eigenvectors, and the underlying worldsheet theory is quantum-exact.
This Hodge–dual minimal area $S[C]$ enters the non-perturbative factor $\exp\left(-\kappa S[C]\right)$ multiplying the perturbative Wilson loop. This factor is an exact zero mode factor satisfying the MM equation in QCD [MMEq79, MM1981NPB, Mig83]. The QCD OPE provides the estimate (12.8) relating our mass scale $\kappa$ to the gluon condensate and QCD mass scale. The primary physical implication is a first-principles geometric proof of confinement, as stated in our Confinement Theorem (Theorem 9.1). This theorem provides an explicit, shape-dependent formula for the effective string tension $\sigma_{\mathrm{eff}}[C]$ (defined as a ratio of $S[C]$ to the Euclidean area $A[C]$). This parameter depends on the loop shape and is bounded from below as $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$. This solution lays the groundwork for a quantitative theory of the high-level hadron spectrum.
Acknowledgements.—The author thanks colleagues at the Institute for Advanced Study for inspiring discussions.