### v3

# Geometric Solution of the Loop Equation and Quark Confinement in QCD

Abstract
We present an exact analytic confining area law for pure $\mathrm{SU}(N_{c})$ Yang–Mills theory, satisfying the hierarchy of multiloop equations for arbitrary regular loops and an arbitrary number of colors $N_{c}$. The solution is constructed using a quaternionic Hodge-dual minimal surface in a 16-dimensional auxiliary space. Unlike random surfaces, this construction defines an effectively topological string theory in which the bulk dynamics reduces to the area of a conformal map of the two-dimensional disk, with all nontrivial dynamics encoded in the boundary loop.
The expectation value of a Wilson loop factorizes into a confining dressing factor and a perturbative term,
$W[C]=W_{\text{pert}}[C]\,e^{-\kappa S[C]}.$
We show that the geometric factor $e^{-\kappa S[C]}$ is an exact multiplicative zero mode of the loop-space diffusion operator which satisfies the full MM multiloop hierarchy for any finite $N_{c}$. The same confining area law applies equally to the planar limit and to the physical case $N_{c}=3$. The functional $S[C]$ obeys the inequality $S[C]\geq\sqrt{2}\,A[C]$ (where $A[C]$ is the Euclidean minimal area), which provides a sufficient condition for confinement.
The area law is given explicitly by the spectral formula:
$S[C]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$
in terms of the two largest eigenvalues of the Douglas–Gram matrix of the loop. We demonstrate that this functional is an exact additive zero mode of the loop diffusion operator, protected algebraically by the loop-space Bianchi identity, and that it is globally defined for all regular loops $C\in H^{1/2}(\mathbb{S}^{1},\mathbb{R}^{4})$. Matching this confining factor with the gluon condensate via the QCD operator product expansion (OPE) determines the physical string tension, $\kappa\sim\Lambda_{\text{QCD}}^{2}$ up to a universal numerical factor. This framework yields an analytic, first-principles derivation of an area law for Wilson loops in QCD and provides new tools for computing the high-lying meson spectrum.
[…]
[aff1]organization=Institute for Advanced Study, city=Princeton, country=USA
1 Introduction
Since the pioneering work of Ken Wilson [Wilson1974Confinement], understanding confinement in Yang–Mills theory remains one of the central challenges of modern physics. In the Wilson loop formulation, confinement is expected to emerge as an area law for large loops. However, in the usual local Lagrangian description, the geometric structure of the loop is largely hidden. An appealing alternative is to reformulate the dynamics directly in loop space, where the MM equations Ref. [MM1981NPB, Mig83] provide a nonlocal, but exact, description of pure $\mathrm{SU}(N_{c})$ Yang–Mills theory in terms of Wilson loop functionals.
In this work, we adopt the loop–space calculus developed in Ref. [migdal2025SQYMflow] to recast the confinement problem in purely geometric terms and to solve the resulting loop equation analytically. In Ref. [migdal2025SQYMflow] the general Euler–Lagrange (EL) equations for a self–dual/anti–self–dual (SD/ASD) Hodge–dual area functional were derived in an auxiliary quaternionic target space, and an exact solution was found for the restricted class of planar loops. The corresponding Hodge–dual minimal surface yields an area law that exactly solves the loop equation for the Yang–Mills gradient flow, but only for loops whose Douglas–Gram matrix has rank two. Moreover, the proof of the planar solution in [migdal2025SQYMflow] contained an error ($Y^{A}_{\mu}=0$), which we correct here. We validate the claim of that paper for planar loops, and generalize it for arbitrary loops.
The first main result of the present paper is a complete holomorphic construction of the quaternionic Hodge–dual minimal surface for arbitrary regular loops $C\in H^{1/2}(\mathbb{S}^{1},\mathbb{R}^{4})$ (equivalently, loops with Fourier coefficients decreasing faster than $1/n$). We show that the minimized Hodge–dual area can be written in a simple spectral form:
$S[C]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$
where $\lambda_{i}$ are the ordered eigenvalues of the Douglas–Gram matrix Ref. [Douglas1931] associated with the loop. This geometric functional is defined by a min–max–min principle (over bulk embeddings, internal isometries, and boundary reparametrizations), is additive under canonical splitting at self–intersections, and has an SD/ASD area derivative protected by the loop–space Bianchi identity. As a consequence, $S[C]$ is an exact zero mode of the loop–space diffusion operator, and the multiplicative dressing
$W[C]\;\longrightarrow\;W[C]\;e^{-\kappa S[C]},$
preserves the loop equation for the Yang–Mills gradient flow for any value of the parameter $\kappa$.
The second main result is that this geometric dressing actually extends beyond the planar approximation: the same functional $S[C]$ defines an exact multiplicative zero mode of the full MM multiloop hierarchy for an arbitrary number of colors $N_{c}$. The key ingredients here are the strict additivity of $S[C]$ at all intersection points and the fact that the same loop operator acts on every member of the hierarchy. When we apply the dressing
$W_{n}(C_{1},\dots,C_{n})\;=\;$
$W_{n,\mathrm{pert}}(C_{1},\dots,C_{n})\,\exp\left(-\kappa\sum_{i=1}^{n}S[C_{i}]\right)$ (1.1)
to the chain of multiloop equations, the geometric factor factors out of each term, cancels between the two sides of the equations, and leaves the entire hierarchy invariant for any finite $N_{c}$. Thus, the confining area law originally obtained in the planar limit is in fact a universal feature of the pure Yang-Mills vacuum, equally valid for the physical case $N_{c}=3$.
A further consequence is a rigorous confinement theorem. Comparing our spectral functional with the Euclidean minimal area $A[C]$ (or, equivalently, with the trace of the Douglas–Gram matrix), we prove the inequality
$S[C]\;\geq\;\sqrt{2}\,A[C],$
which provides a universal lower bound on the effective string tension associated with any regular loop. Finally, by matching the small–loop expansion of the dressed Wilson loop to the operator product expansion (OPE) in QCD, we relate the parameter $\kappa$ to the gluon condensate and the QCD scale, $\kappa\sim\Lambda_{\mathrm{QCD}}^{2}$ up to a universal numerical factor. In this way, the geometric Hodge–dual minimal surface furnishes an exact confining area–law solution of the loop equations in pure $\mathrm{SU}(N_{c})$ Yang–Mills theory and a quantitative bridge to the computation of the hadron spectrum.
2 Basic definitions and duality equations
Our main goal is to minimize the self–dual (or anti–self–dual) area functional
$S_{\chi}[C;X]$ $=\min_{X\in\mathcal{M}_{\rm SD/ASD}}\int_{\mathcal{D}}\!\mathrm{d}^{2}z\,\sqrt{\tfrac{1}{2}\,\Sigma_{\mu\nu}\Sigma_{\mu\nu}},$ (2.1)
where $\Sigma_{\mu\nu}$ is the two–form constructed from the embedding $X_{\mu}(z,\bar{z})$ and defined below. The minimization is taken over embeddings $X_{\mu}(z,\bar{z})\in\mathbb{R}^{4}\!\otimes\!\mathbb{R}^{4}$ satisfying the self–duality (SD) or anti–self-duality (ASD) constraint $\ast\Sigma_{\mu\nu}=\pm\,\Sigma_{\mu\nu}$.
Quaternionic matrix structure. Let $M_{2}(\mathbb{C})$ carry the Hilbert–Schmidt inner product $\langle M,N\rangle_{\mathrm{HS}}=\tfrac{1}{2}\,\operatorname{tr}(MN^{\dagger})$ for $M,N\in M_{2}(\mathbb{C})$. Choose the orthonormal basis $T^{4}=I_{2}$, $T^{i}=\mathrm{i}\sigma_{i}$ ($i=1,2,3$), so that $\operatorname{tr}(T^{A}T^{B\dagger})=2\,\delta^{AB}$. For $X_{\mu}(z,\bar{z})\in M_{2}(\mathbb{C})$, $\mu=1,\dots,4$, define $X^{A}_{\mu}=\tfrac{1}{2}\,\operatorname{tr}(X_{\mu}T^{A\dagger})$, $X_{\mu}=\sum_{A=1}^{4}X^{A}_{\mu}\,T^{A}$.
Complexification of Cartesian components. $X^{A}_{z_{1}}=X^{A}_{1}+\mathrm{i}X^{A}_{2}$, $X^{A}_{\bar{z}_{1}}=X^{A}_{1}-\mathrm{i}X^{A}_{2}$; $X^{A}_{z_{2}}=X^{A}_{3}+\mathrm{i}X^{A}_{4}$, $X^{A}_{\bar{z}_{2}}=X^{A}_{3}-\mathrm{i}X^{A}_{4}$.
Two–form and Hodge decomposition. Define (with Einstein summation over $A$ implied)
$\Sigma_{\mu\nu}=\mathrm{i}\left(\partial_{z}X^{A}_{\mu}\,\partial_{\bar{z}}X^{A}_{\nu}-\partial_{\bar{z}}X^{A}_{\mu}\,\partial_{z}X^{A}_{\nu}\right).$ (2.2)
With $(\ast\Sigma)_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\rho\sigma}\Sigma_{\rho\sigma}$, the SD/ASD parts are $\Sigma^{(\pm)}_{\mu\nu}=\tfrac{1}{2}(\Sigma_{\mu\nu}\pm\ast\Sigma_{\mu\nu})$, satisfying $\ast\Sigma^{(\pm)}_{\mu\nu}=\pm\Sigma^{(\pm)}_{\mu\nu}$. The Euler-Lagrange variational equations for this functional were written down in [migdal2025SQYMflow]. These nonlinear differential equations were solved in that paper only for a special case of a planar boundary loop $C$. In this paper, we generalize that solution for an arbitrary loop.
Self-dual area derivative. There is an internal $O(4)$ gauge invariance of our area element in addition to a physical $O(4)$ invariance of the problem in $\mathbb{R}^{4}$. We use that extra degree of freedom to provide the self-duality (SD or ASD) of the area derivative [migdal2025SQYMflow]. The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
$\frac{\delta W}{\delta\sigma_{\mu\nu}(\theta)}=\frac{\delta^{2}W}{\delta C^{\prime}_{\mu}(\theta-0)\delta C^{\prime}_{\nu}(\theta+0)}-\mu\leftrightarrow\nu$ (2.3)
The loop equation for the Yang–Mills gradient flow was shown in [migdal2025SQYMflow] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
$\mathcal{L}(S)=0;$ (2.4)
$\mathcal{L}=\oint_{\theta}C^{\prime}_{\nu}(\theta)\left(\frac{\delta}{\delta C^{\prime}_{\mu}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\mu}(\theta-0)}\right)\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}$ (2.5)
The basic observation made in that paper was that there is an SD solution similar to the Yang–Mills multi-instanton. The loop space diffusion equation (2.5) is identically satisfied for any functional $S[C]$ with SD or ASD area derivative
$\ast\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)}=\pm\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)};$ (2.6)
$e_{\alpha\mu\nu\lambda}\left(\frac{\delta}{\delta C^{\prime}_{\alpha}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\alpha}(\theta-0)}\right)\frac{\delta S}{\delta\sigma_{\mu\nu}(\theta)}\equiv 0;$ (2.7)
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [migdal2025SQYMflow] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents the SD or ASD tensor. This choice would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
3 Geometric Definition of the Physical Area
The construction of the action $S_{\chi}$ in Sec. 4 depends on a specific choice of boundary conditions (BCs), which introduces a rotational ambiguity. The resulting functional is not $O(4)$ invariant. We must therefore provide a geometric definition of the physical area $S_{\mathrm{phys}}[C]$ that is fully invariant.
We define $S_{\mathrm{phys}}[C]$ as the solution of the minimax problem: first, we minimize the value of the Hodge-dual action $S[X]$ (2.1), by solving the Euler-Lagrange (E-L) equations for $S$ with specific boundary conditions $X=Q.C$ using a normalized matrix $Q$ compatible with the self-duality of the area derivative (see below). Then, we maximize this minimal value over the full group of allowed isometries of $Q$ (transformations that leave the E-L equations invariant).
Case 1: Simple (non-intersecting) loops
For a single, non-intersecting loop $C$, the E-L equations and local Dirichlet BCs $X=QC$ are global. The only relevant isometry that preserves the problem is a global $O(4)$ rotation, $R$. The rotation of $Q$ is equivalent to the opposite rotation of $C$ by the invariance of our area integral. The set of compatible solutions is thus $\{X[RC]\mid R\in O(4)\}$. The definition of the area becomes:
$S_{\mathrm{phys}}[C]:=\max_{R\in O(4)}S[X[RC]]$ (3.1)
This is precisely the definition that will be evaluated in Sec. 5, which yields the $O(4)$-invariant spectral law.
Case 2: Composite (self-intersecting) loops
For a loop that is a composite of two (or more) separated or self-intersecting sub-loops, $C=C_{1}\uplus C_{2}$, the problem changes. The E-L equations and the Dirichlet BCs are local. The problem of finding the minimal surface $X$ for the boundary $C$ therefore decouples into two independent minimal surface problems: $X_{1}$ on domain $\mathcal{D}_{1}$ bounded by $Q_{1}C_{1}$, and $X_{2}$ on domain $\mathcal{D}_{2}$ bounded by $Q_{2}C_{2}$. The total action for the solution is the sum $S[X]=S[X_{1}]+S[X_{2}]$.
Because the problem decouples, the isometry group of the total solution is larger. It is not the "diagonal" $O(4)$ group (which would force $R_{1}=R_{2}$), but the full product group $O(4)_{1}\times O(4)_{2}$. A transformation $(R_{1},R_{2})$ with $R_{1}\neq R_{2}$ is a valid isometry of the decoupled E-L equations.
To find the true maximum over "all compatible isometries" (per our definition), we must maximize over this larger group. This defines the area for a composite loop as $S_{\mathrm{phys}}[C_{1}\uplus C_{2}]=\max_{R_{1}\in O(4)}S[X_{1}[R_{1}C_{1}]]+\max_{R_{2}\in O(4)}S[X_{2}[R_{2}C_{2}]]=S_{\mathrm{phys}}[C_{1}]+S_{\mathrm{phys}}[C_{2}]$.
Finally, the maximized minimal projection area depends on the parametrization of the bounding curve, unlike the Wilson loop, which is reparametrization-invariant by definition. This last ambiguity is eliminated in the same way as in the Euclidean minimal surface: by minimizing the resulting functional over the group of orientation-preserving diffeomorphisms of the circle, $C(\theta)\to C(\phi(\theta))$ with $\phi^{\prime}>0$. Combining these steps, the full definition for a loop decomposable into sub-loops $C_{i}$ is:
$S[C]=\sum_{C_{i}\in C}$
$\min_{\phi\in\mathrm{Diff}^{+}(S^{1})}\max_{R\in O(4)}\min_{\begin{subarray}{c}X:\mathbb{D}_{i}\to\mathbb{R}^{4}\otimes\mathbb{R}^{4}\\ X|_{\partial\mathbb{D}_{i}}=QR(C_{i}\circ\phi)\end{subarray}}\int_{\mathbb{D}_{i}}\lVert dX^{A}\wedge dX^{A}\rVert$ (3.2)
4 A Self-Dual Holomorphic Ansatz
Holomorphic map.
We work in the target complex basis $z_{1}=x_{1}+\mathrm{i}x_{2}$, $z_{2}=x_{3}+\mathrm{i}x_{4}$. Let $U^{A}_{z_{a}}(z)$ be the holomorphic $(1,0)$ components of the embedding $X^{A}_{\mu}(z,\bar{z})=U^{A}_{\mu}(z)+\overline{U^{A}_{\mu}(z)}$. We also define $V^{A}_{z_{a}}:=\partial_{z}U^{A}_{z_{a}}$. As we shall see, the EL equations for the minimal Hodge-dual surface are identically satisfied by virtue of the Cauchy-Riemann equations for the holomorphic functions $U^{A}_{\mu}(z)$ combined with the SD property of the area element.
Lemma 4.1 (Normalization).
With the canonical SD $(1,1)$ form $B:=\mathrm{i}\big(dz_{1}\!\wedge d\bar{z}_{1}+dz_{2}\!\wedge d\bar{z}_{2}\big)$, the Euclidean Hodge star gives $\sqrt{\tfrac{1}{2}\,B_{\mu\nu}B_{\mu\nu}}=2\sqrt{2}$. Consequently, the area density of any SD surface $\Sigma=a(z)B$ is
$\sqrt{\tfrac{1}{2}\,\Sigma^{2}}\;$ = $=\;2\sqrt{2}\;a(z)$
$=\;\sqrt{2}\sum_{A}\big(|V^{A}_{z_{1}}|^{2}+|V^{A}_{z_{2}}|^{2}\big),$ (4.1)
, where the second line follows from $\Sigma_{z_{1}\bar{z}_{1}}=\Sigma_{z_{2}\bar{z}_{2}}=\mathrm{i}a(z)$ and the definition $\Sigma_{z_{a}\bar{z}_{b}}=\mathrm{i}\sum_{A}V^{A}_{z_{a}}\overline{V^{A}_{z_{b}}}$.
Proof.
In the real coframe $e^{1}=dx_{1},\dots,e^{4}=dx_{4}$, $\mathrm{i}dz_{1}\!\wedge d\bar{z}_{1}=2\,e^{1}\!\wedge e^{2}$ and $\mathrm{i}dz_{2}\!\wedge d\bar{z}_{2}=2\,e^{3}\!\wedge e^{4}$, so $B=2(e^{1}\!\wedge e^{2}+e^{3}\!\wedge e^{4})$. Then $\frac{1}{2}B^{2}=8$ and the identities follow. ∎
Two–form and Self-Duality.
The two-form is $\Sigma=\mathrm{d}U\wedge\mathrm{d}\overline{U}$. As $U$ is holomorphic, $\Sigma$ is of pure $(1,1)$ type. The self-duality condition $\ast\Sigma=\Sigma$ requires $\Sigma_{z_{1}\bar{z}_{1}}=\Sigma_{z_{2}\bar{z}_{2}}$. Using the component definition $\Sigma_{z_{a}\bar{z}_{b}}=\mathrm{i}\sum_{A}V^{A}_{z_{a}}\overline{V^{A}_{z_{b}}}$ from Lemma 4.1, this condition implies the explicit constraint:
$\sum_{A}\big(|V^{A}_{z_{1}}|^{2}\big)=\sum_{A}\big(|V^{A}_{z_{2}}|^{2}\big).$ (4.2)
The area density is given by Eq. (4.1).
An Explicit SD Ansatz.
We propose a minimal ansatz that satisfies the self-duality condition (4.2) using two independent holomorphic functions $f_{1}(z)$ and $f_{2}(z)$. Let the only nonzero components of $U$ be:
$U^{4}_{z_{1}}=f_{1}(z),\hskip 18.49988ptU^{3}_{z_{2}}=f_{1}(z),$
$U^{2}_{z_{1}}=f_{2}(z),\hskip 18.49988ptU^{1}_{z_{2}}=f_{2}(z).$ (4.3)
This choice manifestly satisfies (4.2), as $\sum_{A}|V^{A}_{z_{1}}|^{2}=|f_{1}^{\prime}|^{2}+|f_{2}^{\prime}|^{2}$ and $\sum_{A}|V^{A}_{z_{2}}|^{2}=|f_{2}^{\prime}|^{2}+|f_{1}^{\prime}|^{2}$.
Boundary Conditions and Explicit Projector.
The action functional (4.1) for this ansatz becomes a Dirichlet functional
$S_{\chi}=2\sqrt{2}\int_{\mathcal{D}}\big(|f_{1}^{\prime}|^{2}+|f_{2}^{\prime}|^{2}\big)\mathrm{d}^{2}z.$ (4.4)
The EL equations of this functional are satisfied by the Cauchy-Riemann equations for $f_{1},f_{2}$, so that this functional reduces to the boundary data. By the Dirichlet identity, $S_{\chi}=2\sqrt{2}\sum_{n\geq 1}n\big(|\hat{f}_{1,n}|^{2}+|\hat{f}_{2,n}|^{2}\big)$. We now fix the boundary conditions by relating $f_{1},f_{2}$ to the boundary loop $C_{\mu}$ in a rotated basis:
$\Re f_{1}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{1}(\theta)+C_{2}(\theta)\big),$
$\Re f_{2}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{3}(\theta)+C_{4}(\theta)\big).$ (4.5)
Substituting the Fourier modes $\hat{f}_{1,n}=\frac{1}{\sqrt{2}}(\hat{C}_{n,1}+\hat{C}_{n,2})$ and $\hat{f}_{2,n}=\frac{1}{\sqrt{2}}(\hat{C}_{n,3}+\hat{C}_{n,4})$ into (4.4):
$S_{\chi}$ $=2\sqrt{2}\sum_{n\geq 1}n\left(\Big|\frac{\hat{C}_{n,1}+\hat{C}_{n,2}}{\sqrt{2}}\Big|^{2}+\Big|\frac{\hat{C}_{n,3}+\hat{C}_{n,4}}{\sqrt{2}}\Big|^{2}\right)$
$=\sqrt{2}\sum_{n\geq 1}n\left(|\hat{C}_{n,1}+\hat{C}_{n,2}|^{2}+|\hat{C}_{n,3}+\hat{C}_{n,4}|^{2}\right).$ (4.6)
Introducing the Douglas-Gram matrix $G_{\mu\nu}:=\sum_{n\geq 1}n\,\hat{C}_{n,\mu}\,\overline{\hat{C}_{n,\nu}}$, this action is a quadratic form $S_{\chi}=2\sqrt{2}\,\operatorname{tr}(PG)$, where $P$ is the explicit rank-2 projector $P=\frac{1}{2}\left(\begin{smallmatrix}1&1&0&0\\ 1&1&0&0\\ 0&0&1&1\\ 0&0&1&1\end{smallmatrix}\right)$.
Lemma 4.2 (Projector Property).
The matrix $P$ defined by (4) is a rank-2 orthogonal projector ($P\in\operatorname{Gr}(2,4)$). This is verified by direct computation: $P$ is symmetric, $P^{2}=P$, and $\operatorname{tr}P=2$, so $\operatorname{rank}(P)=2$.
Regular loops.
Throughout, we assume the boundary loop $C:S^{1}\to\mathbb{R}^{4}$ is regular in the sense that the spectral expansion converges in the Douglas–Gram matrix, so that the Dirichlet identity leading to (4.6) is valid, and that the Fourier/Hilbert-transform manipulations used in this paper are justified. This corresponds to $C\in H^{\tfrac{1}{2}}(S^{1},\mathbb{R}^{4})$
5 Invariant Area and the Spectral Law
Rotationally invariant area.
The true, physical $O(4)$-invariant area $S_{\mathrm{phys}}$ was defined as the maximum area that can be generated by our functional with all possible rotations $R$ of the loop $C$. For this holomorphic solution
$S_{\mathrm{phys}}[\mathcal{C}]$ $\;:=\;\max_{R\in O(4)}S_{\chi}[RC;P]$
$\;=\;\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}\!\left(P\,(RGR^{\top})\right).$ (5.1)
Let $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$ be the eigenvalues of $G$. The maximum in (5.1) is evaluated by first applying the cyclic property of the trace, which gives $\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}((R^{T}PR)G)$. We can define a new projector $P^{\prime}=R^{T}PR$. Since $P$ is a fixed rank-2 projector, $P^{\prime}$ is also a rank-2 projector. The group $O(4)$ acts transitively on the space of all 2-planes, $\operatorname{Gr}(2,4)$. This transitivity means that by maximizing over all $R\in O(4)$, the projector $P^{\prime}$ sweeps through the entire set of all possible rank-2 projectors. Therefore, our maximization is mathematically identical to $2\sqrt{2}\max_{P^{\prime}\in\operatorname{Gr}(2,4)}\operatorname{tr}(P^{\prime}G)$. By the Ky Fan theorem [Horn:2012:MatrixAnalysis:2nd], this maximum is precisely the sum of the two largest eigenvalues, $\lambda_{3}+\lambda_{4}$.
Outcome and the spectral law. Crucially, the result of this maximization is $S_{\mathrm{phys}}[\mathcal{C}]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})$. This value is a universal spectral invariant, independent of the specific rank-2 projector $P$ we started with.
$S_{\mathrm{phys}}[\mathcal{C}]\;=\;2\sqrt{2}\,(\lambda_{3}+\lambda_{4}).$ (5.2)
This calculation confirms that our physical definition of the invariant area (5.1) leads directly to the $O(4)$-invariant spectral law. For planar loops ($\operatorname{rank}G=2$), $\lambda_{1}=\lambda_{2}=0$, and this correctly specializes to $S_{\mathrm{phys}}[\mathcal{C}]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})=2\sqrt{2}\,\mathcal{D}[C]$, reproducing the claim of Theorem 1 of [migdal2025SQYMflow]. The optimal projector $P_{\ast}=v_{3}v_{3}^{\top}+v_{4}v_{4}^{\top}$ (where $v_{i}$ are eigenvectors of $G$) is the one that achieves this maximum.
5.1 Projector Alignment
The physical area $S_{phys}$ is realized by maximizing the base functional $S_{\chi}$ over the global $\mathrm{O}(4)$ rotation group. The maximum is achieved when the boundary projector aligns with the principal axes of the loop’s geometry:
$S_{phys}=\max_{\mathbf{R}\in\mathrm{O}(4)}2\sqrt{2}\operatorname{tr}\left(\mathbf{P}\mathbf{R}\mathbf{G}\mathbf{R}^{\top}\right).$ (5.3)
The initial fixed projector $\mathbf{P}$ is defined by the orthonormal basis $U=(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3},\mathbf{u}_{4})$ associated with the Hodge parity $\chi=\pm 1$. Explicitly, the basis vectors are:
$\mathbf{u}_{1}$ $=\tfrac{1}{\sqrt{2}}(\mathbf{e}_{1}-\mathbf{e}_{2}),$ $\mathbf{u}_{2}$ $=\tfrac{1}{\sqrt{2}}(\mathbf{e}_{3}-\chi\mathbf{e}_{4}),$
$\mathbf{u}_{3}$ $=\tfrac{1}{\sqrt{2}}(\mathbf{e}_{1}+\mathbf{e}_{2}),$ $\mathbf{u}_{4}$ $=\tfrac{1}{\sqrt{2}}(\mathbf{e}_{3}+\chi\mathbf{e}_{4}).$ (5.4)
In this basis, the projector takes the diagonal form:
$\mathbf{P}=U\operatorname{diag}(0,0,1,1)U^{\top}.$ (5.5)
Similarly, the Douglas-Gram matrix $\mathbf{G}$ is diagonalized by the orthogonal matrix $V$ formed by its eigenvectors $\{\mathbf{v}_{i}\}$ sorted by the magnitude of their eigenvalues $\lambda_{i}$:
$\mathbf{G}=V\operatorname{diag}(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})V^{\top}.$ (5.6)
The optimal rotation $\mathbf{R}_{\ast}$ is the orthogonal transformation that maps the eigenbasis of $\mathbf{G}$ directly onto the basis of $\mathbf{P}$. This rotation matrix is given by the ratio of the two basis matrices:
$\mathbf{R}_{\ast}=UV^{\top}.$ (5.7)
Substituting this rotation into the trace functional yields the maximal value directly as the dot product of the ordered spectra:
$\operatorname{tr}(\mathbf{P}\mathbf{R}_{\ast}\mathbf{G}\mathbf{R}_{\ast}^{\top})$
$=\operatorname{tr}\left((U\Lambda_{P}U^{\top})(UV^{\top})(V\Lambda_{G}V^{\top})(VU^{\top})\right)$
$=\operatorname{tr}\left(\Lambda_{P}\Lambda_{G}\right)=\lambda_{3}+\lambda_{4}.$ (5.8)
6 Self-Duality of the Area Derivative
Theorem 6.1 (Self-Duality of the Physical Action).
The physical action $S_{\mathrm{phys}}[\mathcal{C}]$ defined in (5.2) has a self-dual (SD) area derivative:
$\ast\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}=\chi\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}.$ (6.1)
Proof.
The physical action $S_{\mathrm{phys}}[\mathcal{C}]$ is the envelope (the maximum) of a family of functionals $S_{R}[C]:=S_{\chi}[RC;P]$, indexed by the rotation $R\in O(4)$, where $S_{\chi}[C;P]$ is the base functional from Sec. 4. The area derivative operator $\mathcal{D}:=\frac{\delta}{\delta\sigma}$ and the Hodge star $\ast$ are both $O(4)$-covariant. This covariance implies that if we can prove self-duality for any single member of this family (e.g., the $R=I$ case, $S_{\chi}[C;P]$), then self-duality holds for all members. The derivative of the envelope $S_{\mathrm{phys}}$ at a loop $C$ is, by definition, the derivative of the specific functional $S_{R_{\text{max}}}[C]$ that achieves the maximum at that point. Since all functionals in the family will be shown to have an SD derivative, $S_{\mathrm{phys}}$ must also have an SD derivative. Therefore, the proof reduces to verifying the self-duality for the base functional $S_{\chi}[C;P]$ (for $R=I$), which we do now.
We begin with the generic formula from [migdal2025SQYMflow] for the first variation of the Hodge-dual area
$\delta S=2\int d\theta\delta X^{A}_{\mu}(\theta)\partial_{\theta}X^{A}_{\nu}(\theta)T_{\mu\nu}$ (6.2)
where $T_{\mu\nu}$ is the normalized area element $\Sigma_{\mu\nu}$. With our holomorphic solution, using the constant $B_{\mu\nu}$ tensor and the Dirichlet functional with the Hilbert transform, we find
$\delta S\propto\iint d\theta_{1}d\theta_{2}H(\theta_{1}-\theta_{2})Q(C(\theta_{1}),\delta C^{\prime}(\theta_{2}));$ (6.3)
$Q(A,B)=A_{2}B_{1}-A_{1}B_{2}+A_{3}B_{4}-A_{4}B_{3}$ (6.4)
Now, the first and second variations are
$\frac{\delta S}{\delta C^{\prime}(\theta_{2})}\propto\int d\theta_{1}H(\theta_{1}-\theta_{2})Q(C(\theta_{1}),I_{4});$ (6.5)
$\frac{\delta^{2}S}{\delta C^{\prime}_{[\mu}(\theta-0)\delta C^{\prime}_{\nu]}(\theta)}\propto\oint d\theta^{\prime}H(\theta^{\prime}-\theta)\Theta(\theta^{\prime}-\theta)\hat{Q}_{\mu\nu};$ (6.6)
where $\hat{Q}_{\mu\nu}$ is the ’t Hooft SD or ASD tensor $\hat{Q}_{\mu\nu}=\eta^{3}_{\mu\nu}\text{ or }\bar{\eta}^{3}_{\mu\nu}$. In the second derivative, by the chain rule, we use the standard Volterra derivative [Gelfand:1963:CalculusVariations] of the loop with respect to its tangent, $\delta C(x)/\delta C^{\prime}(y)=\Theta(x-y)$. This tensor $\hat{Q}$ is manifestly self-dual (as $\ast\hat{Q}_{12}=\chi\hat{Q}_{34}$ and $\ast\hat{Q}_{34}=\chi\hat{Q}_{12}$), which proves that the area derivative for $S_{\chi}[C;P]$ is (anti)self-dual. By the covariance argument above, this property extends to $S_{\mathrm{phys}}$.
Note that the numerical factor $\oint d\theta^{\prime}H(\theta^{\prime}-\theta)\Theta(\theta^{\prime}-\theta)\propto\sum_{\text{odd }n>0}1/n$ logarithmically diverges. However, the SD of the area derivative, valid for the tensor factor $Q_{\mu\nu}$, leads to the cancellation of the whole series in the loop space diffusion equation (2.5) (as argued in Sec. 2). Note that this cancellation is valid irrespective of the cutoff at large $n$ in the Fourier series, so that if we fix this cutoff and substitute the area derivative into the loop equation (2.5), we get exactly zero. Thus, we can still claim a finite solution to the loop equation, since the minimal area is finite and its contribution to the loop space diffusion equation is zero, making it a fixed point. ∎
Remark 6.1.
In the computation of the area derivative we use a generic variation of the functional in the full space of embeddings $X^{A}_{\mu}(u,v)$, evaluated on the holomorphic solution after taking the variation. Had we restricted the variation to the holomorphic submanifold $X^{A}_{\mu}(u,v)=\Re\,U^{A}_{\mu}(z)$ (preserving the complex structure), we would have found a vanishing area derivative, since the extremal area is stationary with respect to variations of the moduli. Thus, variations orthogonal to the solution manifold are essential for obtaining a nontrivial area derivative. It is precisely through these orthogonal variations that the boundary condition X=QCX=Q\,C (which carries the Hodge-parity label $\chi$) enters the variation: for $\chi=+1$ the resulting area derivative is self-dual, whereas for $\chi=-1$ it is anti-self-dual, even though the on-shell scalar value $S_{\chi}[C]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})$ is the same in both cases.
A simple analogy is a family of functions $F(x,y;\chi)$ that share a degenerate extremal curve $(x_{0}(t),y_{0}(t))$, such that the extremal value $F(x_{0}(t),y_{0}(t);\chi)$ is independent of $\chi$ for any $t$, while the Hessian components orthogonal to this curve, $(\partial_{\perp})^{2}F$, evaluated on $(x_{0}(t),y_{0}(t))$ still depend on $\chi$. Likewise, $S_{\chi}[C]$ has the same extremal value for both branches, but its area derivative in loop space retains the SD/ASD label. There is therefore no contradiction between the $\chi$-independence of the minimal area and the $\chi$-dependence of its area derivative, which is the antisymmetric two-form part of the second functional derivative of $S[C]$ with respect to the loop tangent $C^{\prime}(\theta)$.
7 The Hodge parity flip
How does our solution change for the negative Hodge parity $\chi=-1$ (ASD instead of SD)? There are many equivalent forms of the ASD solutions, all equivalent after the isometry transformation $O(4)$ rotation of the matrix $Q\Rightarrow QR$ in the Dirichlet BC $X=QC$. The canonical transformation is a T-parity: reflection $X^{A}_{4}\Rightarrow-X^{A}_{4}$ which in complex basis corresponds to $z_{2}\leftrightarrow\bar{z}_{2}$ in our holomorphic ansatz (4.3). After reflection, the only nonzero components will be
$U^{4}_{z_{1}}=f_{1}(z),\hskip 18.49988ptU^{3}_{\bar{z}_{2}}=f_{1}(z),$
$U^{2}_{z_{1}}=f_{2}(z),\hskip 18.49988ptU^{1}_{\bar{z}_{2}}=f_{2}(z).$ (7.1)
As a result, the $(2,0)$ and $(0,2)$ components of the form $\Sigma$ will still be zero, and the $(0,1),(1,0)$ components change. Now the $\Sigma$ form will become $\Sigma=a(z)\tilde{B}$, where $\tilde{B}=e_{1}\wedge e_{2}-e_{4}\wedge e_{4}$ which has negative Hodge duality. The energy density $a(z)$ remains the same, and so does the area functional. The BC (4.5) change in an obvious way
$\Re f_{1}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{1}(\theta)+C_{2}(\theta)\big),$
$\Re f_{2}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{3}(\theta)-C_{4}(\theta)\big).$ (7.2)
which leaves the projection operator $P$ unchanged.
The area derivative in (6.4) will change in an obvious way, following the change of the area element $B\Rightarrow\tilde{B}$
$\delta S\propto\iint d\theta_{1}d\theta_{2}H(\theta_{1}-\theta_{2})\tilde{Q}(C(\theta_{1}),\delta C^{\prime}(\theta_{2}));$ (7.3)
$\tilde{Q}(A,B)=A_{2}B_{1}-A_{1}B_{2}-A_{3}B_{4}+A_{4}B_{3}$ (7.4)
As expected, the scalar value of the area does not depend on the Hodge parity $\chi$, as it is determined solely by the parity-invariant spectrum of the Douglas-Gram matrix. However, the area derivative—the ‘force’ term entering the loop equation—transforms non-trivially, switching from Self-Dual (SD) to Anti-Self-Dual (ASD) under the parity flip. Since the loop diffusion operator $\mathcal{L}$ annihilates both SD and ASD tensor structures independently via the Bianchi identity, both branches $\chi=\pm 1$ represent valid, degenerate zero modes of the equation. To construct a fully Parity-invariant solution for the Wilson loop in QCD, we must treat these degenerate sectors symmetrically. The physical dressing factor involves the sum of both branches in the exponent:
$\exp\left(-\frac{\kappa}{2}\left(S_{\chi=+1}[C]+S_{\chi=-1}[C]\right)\right).$
Since the extremal area values are identical, $S_{+}=S_{-}=S$, this sum restores the discrete symmetry of the theory without altering the spectral form of the confining area law.
8 Reparametrization and factorization for the mulri-loop hierarchy
8.1 Parametric invariance.
As stated below (5.1) and in Sec. 3, the geometric area is defined by maximization over allowed isometries; we additionally take the infimum over orientation–preserving reparametrizations to remove any dependence on the boundary parameter. We therefore adopt the parametrization–invariant form
$S_{\mathrm{phys}}[C]$
$:=\min_{\phi\in\mathrm{Diff}^{+}(S^{1})}\;\max_{R\in O(4)}\,2\sqrt{2}\;\operatorname{tr}\!\Big(P\,R\,G[C\!\circ\!\phi]\,R^{\top}\Big).$ (8.1)
By construction, the choice of parameter is without loss of generality: one may first reparametrize to a convenient standard $0,2\pi]$ and only then perform the $O(4)$ maximization and the outer minimum over $\phi$.
8.2 Canonical split at a self-intersection.
Let $C(\theta_{1})=C(\theta_{2})$ be a transversal self–intersection and define the two oriented subloops $C_{1}=C[\theta_{1},\theta_{2}]$ and $C_{2}=C[\theta_{2},\theta_{1}\!+\!2\pi]$ as in Sec. 3 (Case 2). Using parametric invariance, choose $\phi_{1},\phi_{2}\in\mathrm{Diff}^{+}(S^{1})$ with $\phi_{1}(S^{1})=[\theta_{1},\theta_{2}]$ and $\phi_{2}(S^{1})=[\theta_{2},\theta_{1}\!+\!2\pi]$ and set $\widetilde{C}_{j}(\tau)=C(\phi_{j}(\tau))$ so that each boundary component is on $0,2\pi]$. For a disconnected boundary, the Douglas–Gram series splits,
$G[\widetilde{C}_{1}\!\uplus\!\widetilde{C}_{2}]$ $=G[\widetilde{C}_{1}]+G[\widetilde{C}_{2}],$ (8.2)
and, as explained in Sec. 3 (Case 2), the E–L problems decouple, and the isometry group enlarges to the product $O(4)_{1}\times O(4)_{2}$.
8.3 Additivity of $S_{\mathrm{phys}}$.
With (8.2), the double extremization factorizes:
$S_{\mathrm{phys}}[C_{1}\!\uplus\!C_{2}]:=2\sqrt{2}\,$
$\min_{\phi_{1},\phi_{2}}\;\max_{\begin{subarray}{c}R_{1}\in O(4)\\ R_{2}\in O(4)\end{subarray}}\Big\{\operatorname{tr}\!\big(P\,R_{1}G[C_{1}\!\circ\!\phi_{1}]R_{1}^{\top}\big)$
+ $\hskip 73.07445pt+\,2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{2}G[C_{2}\!\circ\!\phi_{2}]R_{2}^{\top}\big)\Big\}$ (8.3)
$=\min_{\phi_{1}}\max_{R_{1}}2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{1}G[C_{1}\!\circ\!\phi_{1}]R_{1}^{\top}\big)$
$+\;\min_{\phi_{2}}\max_{R_{2}}2\sqrt{2}\,\operatorname{tr}\!\big(P\,R_{2}G[C_{2}\!\circ\!\phi_{2}]R_{2}^{\top}\big)$
$=\;S_{\mathrm{phys}}[C_{1}]+S_{\mathrm{phys}}[C_{2}].$ (8.4)
The first equality is the definition consistent with Sec. 3 (Case 2); The second one uses the product structure of the objective (sum of two independent functionals), which makes the $\max$ over $R_{1},R_{2}$ and the $\min$ over $\phi_{1},\phi_{2}$ separate.
8.4 Compatibility with the MM kernel.
The right–hand side of the MM equation [MMEq79, MM1981NPB] is supported on self–intersections via the bilocal kernel $\int\!\!\int d\theta_{1}d\theta_{2}\,C^{\prime}(\theta_{1})\!\cdot\!C^{\prime}(\theta_{2})\,\delta(C(\theta_{1})-C(\theta_{2}))\,\cdots$. Under $\theta_{a}=\phi_{a}(\tau_{a})$ (with $\phi_{a}^{\prime}\!>\!0$),
$d\theta_{a}\,C^{\prime}(\theta_{a})$ $=d\tau_{a}\,\dot{C}(\tau_{a}),$
$\delta\!\big(C(\theta_{1})-C(\theta_{2})\big)$ $=\delta\!\big(C(\phi_{1}(\tau_{1}))-C(\phi_{2}(\tau_{2}))\big),$ (8.5)
So, the kernel is parameterization-invariant. Therefore, the canonical standardization to $0,2\pi]$ used above is legitimate and (8.4) guaranties that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\circ\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.5) and [Mig83, migdal2025SQYMflow]).
Let us stress that this factorization takes place independently of the splitting points $\theta_{1},\theta_{2}$. Thus, for every $\theta_{1},\theta_{2}$, the product of two factors $\exp\left(-\kappa S[C_{1}\right)$ and $\exp\left(-\kappa S[C_{2}]\right)$ exactly equals the factor $\exp\left(-\kappa S[C]\right)$ on the left side of the MM equation; thus, it can be taken out of the integral, after which it cancels on both sides of the MM equation.
8.5 Confining factor in the multiloop hierarchy
Lemma 8.1 (Finite-$N_{c}$ dressing symmetry).
Let $\{W_{n}\}_{n\geq 1}$ denote the $n$-loop Wilson loop averages in pure $\mathrm{SU}(N_{c})$ Yang–Mills theory, satisfying the MM multiloop hierarchy [Mig83]. Let $S[C]$ be the geometric functional constructed above, with the properties
  1. 1.
    $LS=0$, where $L$ is the loop diffusion operator (2.5);
  2. 2.
    strict additivity at intersections: whenever a loop $C$ self-intersects at $x=y$ into subloops $C_{xy}$ and $C_{yx}$, one has
    $S[C]=S[C_{xy}]+S[C_{yx}],$
    and similarly for intersections between distinct loops.
For any real parameter $\kappa$, define the multiplicatively dressed correlators
$\widetilde{W}_{n}(C_{1},\dots,C_{n})$
$:=W_{n}(C_{1},\dots,C_{n})\,\exp\left(-\kappa\sum_{i=1}^{n}S[C_{i}]\right).$ (8.6)
Then, for any finite $N_{c}$, the family $\{\widetilde{W}_{n}\}$ also satisfies the full MM multiloop hierarchy with the same $N_{c}$.
Proof.
The same loop operator $L$ in (2.5) appears on the left-hand side of the general QCD loop equations [Mig83] for arbitrary $N_{c}$. The difference between the planar and finite-$N_{c}$ cases lies only in the right-hand side: at a self-intersection, the planar factorization is replaced by a combination of one- and two-loop averages, but the kernel of the bilocal integral is the same.
At a self-intersection point of a single loop $C$, the finite-$N_{c}$ loop equation involves the combination:
$W_{2}[C_{xy},C_{yx}]-\frac{1}{N_{c}^{2}}\,W_{1}[C_{xy}\circ C_{yx}].$
This equation is one link in the chain relating $W_{n}$ to $W_{n\pm 1}$; for finite $N_{c}$ the chain terminates due to the Fierz identity for $\mathrm{SU}(N_{c})$ [Mig83], a fact recently exploited in lattice studies of the loop equation [kazakovZheng2024].
In the general multiloop case, $W_{n}[C^{1},\dots,C^{n}]$, the hierarchy contains integrals over pairs of points $x,y$ lying on two (not necessarily distinct) loops $C^{i},C^{j}$, again with the standard MM kernel. When $x$ and $y$ lie on two distinct loops $C^{i}$ and $C^{j}$, the corresponding term involves a merging of these loops in $W_{n-1}$ at the intersection point,
$\frac{1}{N_{c}^{2}}\,W_{n-1}[\dots,C^{i}_{xy}\circ C^{j}_{yx},\dots]\;-\;W_{n}[\dots,C^{i}_{xy},C^{j}_{yx},\dots].$
Thus, every equation in the hierarchy relates loop functionals evaluated on collections of loops that differ only by local splittings or mergings at intersection points.
Now substitute the dressing ansatz (8.6) into the multiloop hierarchy. Because $\mathcal{L}S=0$ and $\mathcal{L}$ is linear, the loop operator on the left-hand side passes through the exponential factor and acts only on the perturbative part:
$\mathcal{L}\,\widetilde{W}_{n}(C_{1},\dots,C_{n})$
$=\exp\left(-\kappa\sum_{i}S[C_{i}]\right)\,\mathcal{L}\,W_{n,\mathrm{pert}}(C_{1},\dots,C_{n}).$ (8.7)
In particular, the dressed correlators satisfy the same differential operator $\mathcal{L}$ as the undressed ones.
On the right-hand side, consider first a self-intersection of a single loop $C$. After dressing, the corresponding contribution becomes
$\Big(W_{2,\mathrm{pert}}(C_{xy},C_{yx})-\frac{1}{N_{c}^{2}}W_{1,\mathrm{pert}}(C_{xy}\circ C_{yx})\Big)$
$\exp\left(-\kappa(S[C_{xy}]+S[C_{yx}])\right).$ (8.8)
By strict additivity, $S[C]=S[C_{xy}]+S[C_{yx}]$ at the intersection point $x=y$. Thus, the exponential factor in (8.8) is the same as the factor multiplying the left-hand side, and therefore cancels from the equation.
For an intersection of two distinct loops $C^{i}$ and $C^{j}$, the right-hand side contains terms of the schematic form
$\Big(W_{n,\mathrm{pert}}(C^{i}_{xy},C^{j}_{yx})-\frac{1}{N_{c}^{2}}\,W_{n-1,\mathrm{pert}}(C^{i}_{xy}\circ C^{j}_{yx})\Big)$
$\times\exp\left(-\kappa\big(S[C^{i}_{xy}]+S[C^{j}_{yx}]+\sum_{k\neq i,j}S[C^{k}]\big)\right).$ (8.9)
Again, by additivity at the intersection and for disjoint unions,
$S[C^{i}]+S[C^{j}]$
$=S[C^{i}_{xy}]+S[C^{i}_{yx}]+S[C^{j}_{xy}]+S[C^{j}_{yx}]$
$=S[C^{i}_{xy}\uplus C^{j}_{yx}]+S[C^{j}_{xy}\uplus C^{i}_{yx}]$
$=S[C^{i}_{xy}\uplus C^{j}_{yx}\uplus C^{j}_{xy}\uplus C^{i}_{yx}]$ (8.10)
so the sum $\sum_{i}S[C^{i}]$ is invariant under the local splitting/merging operation. Hence, the same exponential factor multiplies every term in the multiloop equation and cancels between the two sides.
Since this reasoning applies to each equation in the finite-$N_{c}$ hierarchy and uses only (i) the universal kernel and operator $\mathcal{L}$ and (ii) the additivity and zero-mode properties of $S[C]$, it follows that the dressed correlators $\widetilde{W}_{n}$ satisfy the same multiloop MM hierarchy for any finite $N_{c}$. ∎
Remark 8.1.
Note that the perimeter of the set of loops is also an additive functional, factorizing over the right side of the MM chain of equations. However, this is not a Stokes-type functional, so it does not process finite area derivatives, and, as a consequence, it cannot be added to the zero mode of the loop operator $\mathcal{L}$. Such terms arise as quark-mass renormalization corrections in the perturbative Wilson-loop factor; they are linearly divergent and depend on the regularization scheme. In any case, these terms do not affect the area law at large loops, which is responsible for quark confinement.
9 Confinement theorem
Theorem 9.1 (Confinement area law).
Let $G$ be the Douglas–Gram matrix of a regular loop $C$ with eigenvalues $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$. As proven in the previous sections, the Euclidean geometric area and the Hodge–dual area satisfy
$A_{\mathrm{geom}}[C]:=\operatorname{tr}G\;=\;\sum_{i=1}^{4}\lambda_{i},\hskip 18.49988ptS[C]\;=\;2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$ (9.1)
$\sigma_{\mathrm{eff}}[C]:=\frac{S[C]}{A_{\mathrm{geom}}[C]}\;=\;2\sqrt{2}\,\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}.$ (9.2)
Moreover,
$\sqrt{2}\ \leq\ \sigma_{\mathrm{eff}}[C]\ \leq\ 2\sqrt{2},$ (9.3)
with equality on the right for planar spectra ($\lambda_{1}=\lambda_{2}=0$), and on the left for isotropic spectra ($\lambda_{1}=\cdots=\lambda_{4}$).
The quantity $\sigma_{\mathrm{eff}}[C]$ (hence $S[C]$) is O(4)O(4)–invariant and chirality–even: if $C\mapsto RC$ with $R\in O(4)$, then $G\mapsto RGR^{\top}$ leaves $\{\lambda_{i}\}$, and thus the theorem, unchanged. Combined with Lemma 8.1, this implies that the multiplicative factor $\exp\left(-\kappa S[C]\right)$ defines an exact confining area law solution of the MM loop equations for pure $\mathrm{SU}(N_{c})$ Yang–Mills theory with an arbitrary number of colors $N_{c}$, not just in the planar limit $N_{c}=\infty$.
Remark 9.1 (Minimization over reparametrization).
Both functionals, Douglas’s $\sum_{i}\lambda_{i}$ and our $\lambda_{3}+\lambda_{4}$, must still be minimized over the reparametrization of the bounding loop, $C(\theta)\mapsto C(f(\theta))$ with $f^{\prime}>0$. Since the minimization of just the two highest eigenvalues, as opposed to that of the trace, is generally attained at a different parametrization, we can only state that
$\sigma_{\mathrm{eff}}[C]=2\sqrt{2}\,\frac{\min(\lambda_{3}+\lambda_{4})}{\min(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})}.$
However, the lower inequality still holds: $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$, because for every parametrization one has $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}\leq 2(\lambda_{3}+\lambda_{4})$. Therefore, the minimum in the denominator is less than or equal to twice the minimum in the numerator, which proves the lower bound quoted in the Confinement Theorem.
10 Topological String with Nonlinear Endpoint Dynamics
Our area functional is equivalent to the extremal value of the string theory action
$S_{0}[f_{i}]$ $=\int_{\mathbb{D}}d^{2}z\,(\nabla f_{i})^{2}+\int_{\mathbb{D}}d^{2}z\,u_{i}\,\partial_{\bar{z}}f_{i},$ (10.1)
where $f_{1,2}(z)$ are the basis functions introduced in Section 4. On the boundary, we have linear boundary conditions of the form:
$S_{\partial}[f_{i},\rho]$ $=\oint d\theta\,\rho_{i}(\theta)\left[\Re f_{i}\left(e^{i\theta}\right)-Q_{i}(C(\theta))\right],$ (10.2)
with $Q_{1}(C)=(C_{1}+C_{2})/\sqrt{2}$ and $Q_{2}(C)=(C_{3}+\chi C_{4})/\sqrt{2}$, where $\chi=\pm 1$ is the Hodge parity. Integrating out the auxiliary fields $f_{i}$ reproduces the Douglas-type boundary action $\mathcal{D}[QC]=\mathrm{Tr}(PG)$.
This system describes a topological field theory: the action reduces entirely to boundary terms due to the imposed Cauchy-Riemann equations. Physically, this action represents the net area of two conformal maps, $f_{1}(\mathbb{D})$ and $f_{2}(\mathbb{D})$, with no remaining bulk degrees of freedom. In string theory language, there are no vibrations of the string body; there are only the nonlinear dynamics of its endpoints determined by the loop shape. In this sense, it is analogous to 2D quantum gravity, where the Einstein-Hilbert action reduces to the Euler characteristic, independent of the bulk geometry.
The self-duality of the area derivative stems from local Hodge duality and holds quantum-mechanically. The quadratic nature of the action ensures that the quantum Schwinger–Dyson equations match the classical Euler-Lagrange relations. Thus, the expectation value $\langle\delta S/\delta\sigma\rangle$ is self-dual, and the quantum string obeys the classical loop equations.
Integrating out $f_{i}$ yields a Gaussian integral:
$Z[C;P]$ $\propto[\det{}^{\prime}(K_{P})]^{-1/2}$
$\times\exp\left(-2\sqrt{2}\,\kappa\mathrm{Tr}(PG)\right),$ (10.3)
where $K_{P}=P\,H\,\partial_{\theta}P$ has a universal spectrum. Hence $\det{}^{\prime}(K_{P})$ is a constant, and the effective action is classical: $\Gamma[C;P]=2\sqrt{2}\,\mathrm{Tr}(PG)+\text{const}$.
According to our min-max-min principle, which is required to preserve the factorization of the product of dressing factors $\exp\left(-\kappa S[C]\right)$ on the right side of the MM equation, we do not integrate over the “moduli space” (the rotation matrix $R$). Instead, we select the optimal rotation that brings the exponent to its maximal value:
$Z[C]$ $\propto\exp\left(-2\sqrt{2}\,\kappa\max_{R\in SO(4)}\mathrm{Tr}(R^{T}PRG)\right)$
$=\exp\left(-2\sqrt{2}\,\kappa(\lambda_{3}+\lambda_{4})\right).$ (10.4)
This prescription corresponds to setting an optimal isometry of the boundary condition rather than averaging over these isometries. In terms of the partition function $Z=\exp\left(-\kappa S[RC]\right)$, our prescription selects the minimal value, rather than performing a saddle point integration around the maximal value of this factor.
Let us stress that we are solving the QCD loop equations rather than quantizing a fundamental string theory. The internal degrees of freedom found in non-critical string theory (such as Polyakov’s Liouville mode) do not apply here. We utilize string theory as a mathematical tool, acting as a computational dual. This approach mirrors the logic of the Eguchi-Kawai reduction [8], where auxiliary variables (quenched momenta) are introduced not as physical degrees of freedom, but as instruments to satisfy the loop equations.
In our picture, the fluctuations of this electric flux are described by gluon graphs in another factor $W_{pert}[C]$. The confining factor $Z[C]$ only reflects the mean properties of the vacuum.
The optimal rotation $R_{\ast}$ described in section 5.1, aligning the SD/ASD projector with the highest plane of the Gram basis, determines the boundary conditions for the holomorphic string:
$\Re f_{i}\left(e^{i\theta}\right)=Q_{i}(R_{*}C(\theta)).$ (10.5)
For the ASD string, the reflected coordinate $C_{4}$ is used in $Q_{i}(RC)$.
11 String Scale $\kappa$ and Gluon Condensate
The string mass scale $\kappa$ is not fixed by the MM equation itself. As we have seen, the confining factor $\exp\left(-\kappa S[C]\right)$ passes through the equation without correlation with the perturbative solution $W_{pert}[C]$. This factor reflects the nonperturbative properties of the QCD vacuum rather than the perturbative fluctuations of the gluon field.
The fundamental nonperturbative property of the QCD vacuum is the gluon condensate [Shifman:1978bx]. We relate the string tension $\kappa$ to this condensate by considering a circular loop $C$ of small radius $r$. According to our solution, $S[C]=2\sqrt{2}\pi r^{2}$. We compute the second area derivative at two opposite points $x_{1}=C(0)$ and $x_{2}=C(\pi)$ on the circle.
At small distances $|x_{12}|=2r$, the Operator Product Expansion (OPE) in asymptotically free QCD is dominated by the leading conformal term in $W_{pert}[C]$:
$\frac{\delta^{2}W_{pert}[C]}{\delta\sigma_{\mu\nu}(1)\delta\sigma_{\lambda\rho}(2)}$ $\propto\langle\mathrm{Tr}F_{\mu\nu}(x_{1})U_{0,\pi}F_{\lambda\rho}(x_{2})U_{\pi,2\pi}\rangle$
$\to\frac{\alpha_{eff}}{4\pi}\frac{I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}}{|x_{12}|^{4}}W_{pert}[C],$ (11.1)
where $I_{\mu\lambda}=\delta_{\mu\lambda}-2\hat{x}_{\mu}\hat{x}_{\lambda}$ with $\hat{x}=x_{12}/|x_{12}|$. We project this perturbative singularity out by contracting with the tensor $\Pi_{\mu\nu\lambda\rho}=\tfrac{1}{2}(\delta_{\mu\lambda}\delta_{\nu\rho}-\delta_{\nu\lambda}\delta_{\mu\rho})$. Simple tensor algebra confirms that the leading term vanishes: $\mathrm{Tr}(\Pi I\otimes I)=0$.
The remaining term is the gluon condensate contribution. This term involves the renormalization factor $Z^{2}$ associated with the two cusps created at the points of differentiation:
$\Pi_{\mu\nu\lambda\rho}\mathrm{Tr}F_{\mu\nu}(1)U_{1,2}F_{\lambda\rho}(2)U_{2,1}\to$
$-\langle(G^{a}_{\mu\nu})^{2}\rangle W_{pert}[0]Z^{2}.$ (11.2)
This constant contribution must be balanced by the second area derivative of our Hodge-dual factor in the ansatz $W=W_{pert}e^{-\kappa S}$. Using the fact that the first derivative $\delta S/\delta\sigma$ vanishes for the circle due to symmetry, we find:
$W_{pert}[C]\frac{\delta^{2}\exp\left(-\kappa S[C]\right)}{\delta\sigma_{\mu\nu}(1)\delta\sigma_{\lambda\rho}(2)}\to$
$-\kappa W_{pert}[0]\Pi_{\mu\nu\lambda\rho}\frac{\delta^{2}S[C]}{\delta\sigma_{\mu\nu}(0)\delta\sigma_{\lambda\rho}(\pi)}.$ (11.3)
Crucially, there is no $Z$ factor in this geometric term; it depends only on the renormalization of the perturbative loop $W_{pert}[0]$ which appears as a common prefactor.
The second area derivative is computed as the second variation of $S=2\sqrt{2}(\lambda_{3}+\lambda_{4})$:
$\frac{\delta^{2}S[C]}{\delta\sigma^{2}}\propto\mathrm{Tr}(\Pi B\otimes B)\sum_{i<j}\frac{\left(\sum_{n>1}\frac{1}{n}\right)^{2}}{\lambda_{j}-\lambda_{i}}.$ (11.4)
The harmonic sum $\sum_{n>1}1/n\to\log n_{max}$ arises from the singular Hilbert kernels in the area derivatives. The result is:
$\frac{\delta^{2}S[C]}{\delta\sigma^{2}}=\frac{a(\log n_{max})^{2}}{r^{2}}.$ (11.5)
The sign matches the gluon condensate term, while the magnitude depends on the radius.
Remark 11.1.
The divergence scaling as $1/r^{2}$ arises from the second area derivative as a direct consequence of the scaling property of our Hodge minimal area. This area is a nonlinear, nonlocal, homogeneous functional of degree two. Every area derivative involves two functional derivatives with respect to $\dot{C}$, lowering the degree by two. Thus, the second area derivative has degree minus two, inevitably scaling as $1/r^{2}$. The logarithmic divergence comes from the high-frequency spectral sum.
The proper physical procedure is to match these terms at the crossover scale. In the matching region, we equate the renormalized OPE result with the geometric derivative:
$Z^{2}\langle(G^{a}_{\mu\nu})^{2}\rangle=\frac{\kappa a(\log n_{max})^{2}}{r^{2}}\exp\left(-2\sqrt{2}\pi r^{2}\kappa\right).$ (11.6)
where the constant $a\sim 1$ is calculable from our solution, and renormalization constant $Z$ for the area derivative operator applied to $W_{pert}[C]$ is calculable by the methods of [KorchemskyRadyushkin1987]. Presumably, the renormalization factor $Z^{2}$ on the LHS compensates for the UV divergence $(\log n_{max})^{2}$ on the RHS, rendering the relation finite and renormalization-group invariant. The Fourier cutoff $n_{max}$ can be related to the UV cutoff $\Lambda$ of QCD by estimating the sides of the polygonal loop with discrete Fourier transform instead of continuum one $\Delta\theta=2\pi/n_{max}\sim 1/(\Lambda|C^{\prime}|)\sim 1/(\Lambda r)$, so that $n_{max}\sim\Lambda r$.
The function $f(z)=ze^{-z}$ (where $z=2\sqrt{2}\pi r^{2}\kappa$) reaches maximum at $z=1$. This limitation implies that no solutions exist outside this parameter region, suggesting the presence of a critical point. We identify this maximum as the natural matching point:
$r_{c}^{2}=\frac{1}{2\sqrt{2}\pi\kappa};$ (11.7)
$\kappa^{2}=\frac{e}{2\sqrt{2}a\pi}\left(\frac{Z}{\log n_{max}}\right)^{2}\left\langle(G^{a}_{\mu\nu})^{2}\right\rangle.$ (11.8)
The factors of $N_{c}$ cancel, consistent with the planar limit. This matching fixes the string tension $\kappa$ in terms of the gluon condensate $\langle G^{2}\rangle$.
Our association of $\kappa$ with the gluon condensate supports the physical picture of condensed electric flux originally conjectured by Nambu [Nambu:1974zg]. One may view the term $\kappa S[C]$ as the action cost of a field configuration with energy density $\langle G^{2}\rangle\sim\Lambda_{QCD}^{4}$ spread over a “pancake” surface of thickness $\Lambda_{QCD}^{-2}$ containing the 44- volume $\Lambda_{QCD}^{-2}S[C]$. This estimate interprets the dressing factor $\exp\left(-\kappa S[C]\right)$ as a statistical probability $\exp\left(-\beta\delta E\right)$ for a large quark loop propagating through the QCD vacuum (the dual Meissner effect).
12 Conclusion
We have presented an exact solution for the quaternionic Hodge–dual minimal surface, a new class of minimal surfaces in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ that can be solved exactly by reduction to free conformal maps. The minimized Hodge–dual area is the nontrivial spectral invariant:
$S[C]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$
where $\{\lambda_{i}\}$ are the eigenvalues of the Douglas–Gram matrix $G$. The two dominant Douglas eigenvectors determine the optimal surface, and the bulk dynamics reduces to a nonlinear boundary functional, depending on the loop geometry. There are no physical degrees of freedom associated with fluctuations of the shape of the surface, unlike ordinary string theory.
The Hodge–dual minimal area $S[C]$ enters the nonperturbative dressing factor $\exp\left(-\kappa S[C]\right)$ multiplying the perturbative Wilson loop. Thanks to the self–dual/anti–self–dual area derivative and the strict additivity of $S[C]$ at intersections, this dressing defines an exact zero mode of the loop–space diffusion operator. By Lemma 8.1, the multiplicative dressing preserves the entire MM multiloop hierarchy for any finite number of colors $N_{c}$, including the physical case $N_{c}=3$. Thus, the same geometric confining area law applies both in the planar limit and at finite $N_{c}$. The QCD OPE analysis provides the estimate (11.8), relating the mass scale $\kappa$ to the gluon condensate and the QCD scale.
The central physical implication is summarized in our Confinement Theorem (Theorem 9.1), which yields an explicit, shape–dependent formula for the effective string tension:
$\sigma_{\mathrm{eff}}[C]:=\frac{S[C]}{A_{\mathrm{geom}}[C]},$
defined as the ratio of the Hodge–dual area $S[C]$ to the Euclidean geometric area $A_{\mathrm{geom}}[C]=\mathrm{tr}\,G$. The effective tension depends on the loop geometry but obeys the universal lower bound $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$, providing a rigorous geometric confinement criterion. Together with the OPE matching, this framework yields a confining area–law solution of the loop equations in pure $\mathrm{SU}(N_{c})$ Yang–Mills theory and lays the basis for a computation of the high–lying meson spectrum.
Acknowledgements.—The author thanks colleagues at the Institute for Advanced Study for inspiring discussions.